### v1

# Prospects of constraining on the polarizations of gravitational waves from binary black holes using space- and ground-based detectors

Abstract
In general relativity, gravitational waves (GWs) exhibit two tensor modes, while alternative theories predict up to six polarization modes. We investigate GW polarization constraints using a model-independent parametrized post-Einsteinian framework with space- and ground-based detectors. Evaluating LISA, Taiji, TianQin, LIGO, Virgo, KAGRA, and Einstein Telescope (ET), we analyze their capabilities and network performance. Taiji provides the best constraints among space-based detectors, with LISA outperforming TianQin. Among ground-based detectors, LIGO excels in vector modes, while ET offers the most comprehensive constraints. In network scenarios, LISA+TJm performs best, and ET surpasses second-generation detector combinations. Multiband observations alleviate scalar mode degeneracies, significantly enhancing ground-based detector performance. Combined space- and ground-based observations yield robust constraints on GW polarization, advancing tests of deviations from general relativity. Our findings highlight the potential of future GW missions to refine our understanding of gravitational physics through precise polarization measurements.
I Introduction
Over the past century, general relativity (GR) has been rigorously tested through numerous solar system experiments and astrophysical observations, with no definitive evidence suggesting deviations from GR [1, 2, 3]. Despite passing all experimental tests, there are indications that GR may need to be extended, particularly due to the challenges posed by dark energy and dark matter in cosmology [4, 5]. Gravitational waves (GWs) offer a new avenue for testing GR, especially in the strong-field regime where GWs are generated by compact objects [6]. One method to test GR is by detecting additional polarizations of GWs. In GR, GWs only have two tensor modes, whereas some alternative gravity theories predict up to six polarization modes [7]. For instance, Brans-Dicke theory predicts an additional scalar mode [2, 8], while $f(R)$ theory and Horndeski theory propose two additional scalar modes [9, 10, 11]. Einstein-Aether theory predicts five polarization modes [12, 13], and some tensor-vector-scalar theories encompass all six polarization modes [14, 15]. Therefore, detecting additional polarizations of GWs can help identify deviations from GR, surpass current experimental limits, and reveal the deeper nature of gravity.
Since LIGO first detected GW [16], nearly a hundred GW events generated by compact binary coalescences (CBCs) have been observed [17, 18, 19]. Testing GR with these GW data has shown that all observed data are consistent with GR so far [20, 21, 22, 23, 24, 25, 26]. With the commencement of the fourth observation run (O4), more GW events are expected to be detected, allowing for further testing of GR. In general, to detect the polarization of transient GWs, more detectors are needed to obtain responses from different directions, thereby enhancing detection capabilities [27]. Ground-based detectors in the hertz frequency band, such as LIGO [28], Virgo [29], and KAGRA [30], collaborate to form a detector network to detect the polarization of GWs. For continuous GWs emitted by distant CBCs, a detector with directional change capability is sufficient [27]. Space-based GW detectors in the millihertz frequency band, such as LISA [31], Taiji [32], and TianQin [33], consist of three spacecraft orbiting the Sun/Earth in a triangular formation, making it possible to detect additional GW polarizations due to the detector’s motion in space [34].
Besides determining whether the detector can theoretically detect additional polarizations of GWs, it is also crucial to ascertain whether the detected GW signal is consistent with the predictions of GR. One approach is to use a waveform model with purely phenomenological parameters to represent possible deviations from GR and to apply constraints to these parameters using observed GW data [6]. Yunes and Pretorius proposed the parameterized post-Einstein (ppE) waveform [35], based on the post-Newtonian (PN) approximation, which is suitable for parameterizing the influence of alternative gravity theories on non-GR polarizations. Most modified gravity theories, such as Brans-Dicke theory, massive gravity, and bimetric theory, can be described using the ppE framework [36]. Although the ppE framework cannot parameterize all possible deviations from GR, it can test for additional polarizations, providing a method to evaluate the performance of detectors in testing GR [37, 38, 39].
Currently, some work utilizes the ppE framework to constrain additional polarizations for testing GR. Narikawa and Tagoshi discussed the potential of advanced ground-based GW detectors, such as LIGO, Virgo, and KAGRA, in detecting general deviations between GWs and the predictions of GR using the ppE framework [40]. Huwyler et al. employed LISA to detect massive black hole binaries (MBHBs) and test GR by representing ppE waveforms with phase corrections only [41]. References [36, 6] discussed the potential of Taiji and TianQin in using MBHBs to test GR and provide constraints on non-GR parameters. For detector networks, Nair et al. focused on the synergistic effect of the Einstein Telescope (ET) and preDECIGO, demonstrating enhanced sensitivity within a specific band [42]. In Ref. [43], Wang and Han used the LISA-Taiji network to test the ppE parameters of deviations from the GR waveform, indicating a significant improvement in the detection of polarization amplitude through joint observations.
Based on our previous work [44, 45, 46], we investigate the constraints from binary black holes (BBHs) for alternative configurations of space- and ground-based detectors on the detection of additional GW polarizations. We employ the ppE framework for model-independent testing in GR by numerically calculating time-domain GW signals that contain all polarizations. Space-based detectors LISA, Taiji, and TianQin are used to observe massive black hole binaries (MBHBs), while ground-based detectors LIGO, Virgo, KAGRA, and the Einstein Telescope (ET) observe stellar-mass binary black holes (SBBHs), providing constraint results under different networks. We also consider the combination of multiband observations. The long-term observation of SBBHs by space-based detectors helps break the response degeneracy of ground-based detectors, thereby improving constraints. Using the Fisher information matrix (FIM), we present the constraint results of several typical mass BBHs as a function of redshift. Additionally, we discuss the potential effects of multimessenger observations on enhancing parameter constraints. Through systematic research, we comprehensively analyze the results of additional GW polarizations constrained by space- and ground-based detectors from multiple perspectives.
This paper is organized as follows. In Sec. II, we introduce time-domain GW signals and ppE parameters for measuring additional polarizations within the ppE framework. In Sec. III, we review the performance parameters and alternative configurations of space- and ground-based detectors, and analyze the response functions corresponding to different modes. In Sec. IV, we explain the typical BBH sources selected in our paper and the method for calculating the signal-to-noise ratio (SNR) and FIM. In Sec. V, we present the constraint results on ppE parameters, including observations from different networks, multiband, and multimessenger observations. Finally, we summarize the results of our research in Sec. VI.
II Gravitational wave signal
Figure 1: Comparison of time-domain GW waveforms under GR and ppE frameworks. The GW waveform of an MBHB with a total mass of $10^{5}\ \mathrm{M_{\odot}}$ is observed by LISA over the year before coalescence. To clearly visualize differences, we set $\alpha_{D}=0.001$, $\alpha_{Q}=96/5$, and $\alpha_{V}=\alpha_{B}=\alpha_{L}=0.15$, ensuring $\mathcal{A}_{V}=\mathcal{A}_{B}=\mathcal{A}_{L}$. (a) displays the specific time-domain GW signal. (b) and (c) show the GW waveforms from 415 days and 30 minutes before coalescence, reflecting the early and late inspiral phases, respectively. (d) illustrates the ratio of the amplitude of additional modes to the tensor modes.
In generalized modified gravity theories, metric perturbations can have up to six independent degrees of freedom, resulting in six different polarization modes: two tensor modes ($+$ and $\times$), two vector modes ($X$ and $Y$) and two scalar modes ($B$ and $L$[47]. For detectors, the observed GW strain can be described as a linear combination of different GW polarizations, expressed as
$h(t)=\sum_{A}F^{A}h_{A}(t),$ (1)
where $A=\{+,\times,X,Y,B,L\}$ represents for the six polarizations, $h_{A}(t)$ is the input signal of GWs, and $F^{A}$ is the angular response function.
The extended ppE framework is used to construct a model-independent test for GR, including all GW polarization modes [37]. The amplitude and phase of GWs can be obtained separately from the measurement of perturbations and energy evolution, and there exist simultaneous quadrupole and dipole radiation. The process of CBC includes three phases: inspiral, merger, and ringdown. Currently, the observed GW events are transient and do not include the early inspiral phase [19]. Moreover, GR has passed all current experimental tests, indicating that only tensor modes generated by quadrupole radiation are present in the non-inspiral phase. The contribution of dipole radiation in the early inspiral phase can be greater than during the merger phase [36, 48]. Therefore, it is reasonable to assume that tensor modes dominate in quadrupole radiation, while vector and scalar modes dominate in dipole radiation. Following Refs. [49, 36, 37, 47], the GW waveform under the ppE framework can be written as
$h_{+}$ $=\mathcal{A}_{T}(1+\cos^{2}\iota)/2\cos(2\Phi+2\Phi_{0}),$ (2)
$h_{\times}$ $=\mathcal{A}_{T}\cos\iota\sin(2\Phi+2\Phi_{0}),$
$h_{X}$ $=\mathcal{A}_{V}\cos\iota\cos(\Phi+\Phi_{0}),$
$h_{Y}$ $=\mathcal{A}_{V}\sin(\Phi+\Phi_{0}),$
$h_{B}$ $=\mathcal{A}_{B}\sin\iota\cos(\Phi+\Phi_{0}),$
$h_{L}$ $=\mathcal{A}_{L}\sin\iota\cos(\Phi+\Phi_{0}),$
with
$\mathcal{A}_{T}$ $=\frac{4}{D_{L}}\left(\frac{G\mathcal{M}}{c^{2}}\right)^{5/3}\left(\frac{\omega}{c}\right)^{2/3},$ (3)
$\mathcal{A}_{V}$ $=\frac{\alpha_{V}}{D_{L}}\left(\frac{G\mathcal{M}}{c^{2}}\right)^{4/3}\left(\frac{\omega}{c}\right)^{1/3},$
$\mathcal{A}_{B}$ $=\frac{\alpha_{B}}{D_{L}}\left(\frac{G\mathcal{M}}{c^{2}}\right)^{4/3}\left(\frac{\omega}{c}\right)^{1/3},$
$\mathcal{A}_{L}$ $=\frac{\alpha_{L}}{D_{L}}\left(\frac{G\mathcal{M}}{c^{2}}\right)^{4/3}\left(\frac{\omega}{c}\right)^{1/3},$
where $\alpha_{V,B,L}$ are the dimensionless ppE parameters, $\mathcal{M}=(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$ is the chirp mass, $m_{1}$ and $m_{2}$ are the masses of BBH, $D_{L}$ is the luminosity distance, $\iota$ is the inclination angle, $\Phi=\int{\omega}\mathrm{d}t$ is the orbital phase, $\omega$ is the orbital angular frequency, $\Phi_{0}$ is the initial orbital phase, $G$ and $c$ are the gravitational constant and the speed of light.
We only consider the dominant modes in different radiations for the amplitude, and the contribution of different radiations to the orbital angular frequency varies. Considering the influence of possible non-dominant modes, the overall evolution of the orbital angular frequency can be described by the contributions of dipole radiation and quadrupole radiation [49, 36]:
$\frac{\mathrm{d}\omega}{\mathrm{d}t}=\alpha_{D}\eta^{2/5}\frac{G\mathcal{M}}{c^{3}}\omega^{3}+\alpha_{Q}\left(\frac{G\mathcal{M}}{c^{3}}\right)^{5/3}\omega^{11/3},$ (4)
where $\alpha_{D}$ and $\alpha_{Q}$ are the ppE parameters that describe the orbital angular frequency contributions of dipole and quadrupole radiation, $\eta=m_{1}m_{2}/M^{2}$ is the symmetric mass ratio, and $M=m_{1}+m_{2}$ is the total mass.
To describe the model-independent GW waveform, we select five dimensionless ppE parameters to constrain parameters in alternative gravity theories. In the case of GR, $\alpha_{D,V,B,L}=0$ and $\alpha_{Q}=96/5$. To ensure our analysis remains independent of specific models, we treat these ppE parameters as independent. Some theories exhibit correlations between these parameters, potentially enhancing parameter constraints [36]. Thus, our results are conservative, and the actual constraints may exceed those presented here.
From Eq. (4), solving the evolution function $\omega(t)$ of the orbital angular frequency over time analytically is challenging. We determine $t(\omega)$ through integration::
$t=t_{0}+\int_{\omega(t_{0})}^{\omega(t)}{\left(\frac{\mathrm{d}\omega}{\mathrm{d}t}\right)^{-1}}\mathrm{d}\omega,$ (5)
which is shown in Refs. [49, 36]. Computational methods, such as the bisection method, iteratively solve for the orbital angular frequency corresponding to a given time point. By solving point by point, we obtain $\omega(t)$. Using the above method, we input the calculated $\omega(t)$ into Eqs. (1)-(3) to derive the final GW signal.
In order to better demonstrate the impact of ppE parameters on GW waveforms significantly, we set large parameter values in Fig. 1. As shown in Fig. 1(b), there is a notable amplitude difference between the ppE waveform and the GR waveform. This discrepancy arises because the GW frequency from dipole radiation equals the orbital frequency of the BBH, whereas quadrupole radiation operates at twice that frequency. This effectively superimposes a waveform with half the frequency onto the original GR waveform, resulting in amplitude modulation at twice the wavelength. Comparing Figs. 1(b) and (c), the disparity between the ppE waveform and the GR waveform is significant when the merger is distant but diminishes as the merger approaches, consistent with our assumption in Eq. (2) to align with current GR testing outcomes. In Fig. 1(d), the ratio of additional modes to tensor modes is visually depicted, showing a gradual decrease over time, causing the ppE waveform to progressively approach the GR waveform. From Eq. (3), it can be seen that the amplitude of the tensor modes varies with the power of $2/3$ of the angular frequency, while the additional modes varies with the power of $1/3$. The rate of amplitude increase for tensor modes is significantly greater than for additional modes, resulting in the ratio $\mathcal{A}_{V}/\mathcal{A}_{T}\propto\omega^{-1/3}$ declining as angular frequency increases.
III Detectors and response
III.1 Space-based detectors
Figure 2: Configuration and location of space- and ground-based detectors (not to scale). (a) and (b) depict alternative orbital configurations of LISA, Taiji-p, Taiji-c, Taiji-m, TianQin I, and TianQin II [44]. The solid arrow on the constellation plane represents the normal direction of the detector constellation plane. (c) shows the geographical locations and azimuth angles of LIGO, Virgo, KAGRA, and ET [50, 51].
Figure 3: The variation of the normal direction of the detector constellation plane during the observation period in the ecliptic $(\lambda,\beta)$ and equatorial $(\alpha,\delta)$ coordinate systems. This paper uses the obliquity of the ecliptic ($\varepsilon_{2000}=23^{\circ}26^{\prime}21^{\prime\prime}$) based on epoch J2000.0 for the conversion between these two coordinate systems [52].
In the 2030s, LISA, Taiji, and TianQin are scheduled to launch triangular constellations consisting of three spacecraft. LISA and Taiji utilize heliocentric orbits, whereas TianQin employs geocentric orbits. Various alternative orbital configurations are available, as detailed in Figs. 2 and 3. LISA and Taiji, including three configurations, have a leading or trailing angle of $20^{\circ}$ relative to Earth, with a $60^{\circ}$ inclination between the constellation plane and the ecliptic plane [31, 32]. Furthermore, the normal direction of the TianQin constellation plane remains fixed, and TianQin follows a “three months on+three months off” observation scheme [33].
Variations in the constellation plane determine how GW sources appear in the detector coordinate system, thereby influencing the response to different polarizations. Additionally, detector sensitivity is directly influenced by the arm length, with LISA, Taiji, and TianQin having arm lengths of $2.5\times 10^{6}$ km, $3\times 10^{6}$ km, and $\sqrt{3}\times 10^{5}$ km, respectively. Concerning the noise in space-based detectors, we account for acceleration noise, displacement noise, and foreground noise. Acceleration noise and displacement noise contribute to the primary power spectral density (PSD), as detailed in Refs. [53, 54, 55]. Foreground noise originates from galactic binaries within the Milky Way galaxy, creating a peak in the frequency band of approximately 0.3-3 mHz, as demonstrated in our previous work [44].
III.2 Ground-based detectors
In contrast to the planned space-based detectors, the ground-based detectors are currently operational, primarily comprising the LVK collaboration, which includes LIGO, Virgo and KAGRA. These are all L-shaped second-generation detectors with arm lengths of 4 km (LIGO), 3 km (Virgo), and 3 km (KAGRA), respectively [28, 29, 30]. In this paper, we consider LIGO to include three detectors: LIGO Hanford (H1) and LIGO Livingston (L1), which are operational, and the planned LIGO India (I1). Furthermore, we consider the third-generation detector ET, which features a triangular shape composed of three 10 km arms [56].
As the Earth rotates, ground-based detectors scan different sky regions, as illustrated in Fig. 3. Due to the obliquity of the ecliptic, the sky regions scanned by space- and ground-based detectors is not parallel. Ground-based detectors, unlike their space-based counterparts, are affected by various types of noise, including quantum noise, seismic noise, gravity-gradient noise, thermal noise, and others. For our study, we utilize the design performance specifications of these detectors, and the corresponding PSD can be referenced in LIGO Document T1500293 and Refs. [57, 58].
III.3 Response function
Figure 4: Relationship between the detector coordinate and the GW coordinate [45].
Figure 5: Angular response functions for the combined tensor mode $F^{T}$, combined vector mode $F^{V}$, breathing mode $F^{B}$, and longitudinal mode $F^{L}$ at LISA detector coordinates $(\phi_{d},\theta_{d})$. The left panel shows the response functions for GWs at 1 mHz, where the breathing and longitudinal modes are almost identical. The right panel displays the response at twice the transfer frequency, $2f_{*}=38\ \mathrm{mHz}$, where the degeneracy between the breathing and longitudinal modes is broken at higher frequencies.
A dual-arm Michelson interferometer detects GWs by measuring the relative change in the length of its two arms. We describe the detector’s response to GWs using the method outlined in Refs. [45, 27]. As shown in Fig. 4, the detector coordinates are constructed using orthogonal unit vectors $\hat{x},\hat{y},\hat{z}$}, and the GW coordinates are constructed using $\hat{p},\hat{q},\hat{w}$}. Here, $\hat{w}$ represents the propagation direction of the GW, and the unit vector $\hat{\Omega}=-\hat{w}$ represents the position of the GW source. In a triangular detector, the angle $\gamma$ between the two arms is $60^{\circ}$, while in an L-shaped detector, it is $90^{\circ}$. For GWs, an additional rotational degree of freedom can be fixed by specifying the polarization angle $\psi$, ultimately using $\hat{m},\hat{n},\hat{w}$} to describe the GW. The angular response function $F^{A}$ in Eq. (1) can be expressed as
$F^{A}=D^{ij}e_{ij}^{A},$ (6)
where the polarization tensor $e_{ij}^{A}$ is described using the orthogonal unit vectors mentioned above [27, 59]:
$e_{ij}^{+}$ $=\hat{m}_{i}\hat{m}_{j}-\hat{n}_{i}\hat{n}_{j},$ $e_{ij}^{\times}$ $=\hat{m}_{i}\hat{n}_{j}+\hat{n}_{i}\hat{m}_{j},$ (7)
$e_{ij}^{X}$ $=\hat{m}_{i}\hat{w}_{j}+\hat{w}_{i}\hat{m}_{j},$ $e_{ij}^{Y}$ $=\hat{n}_{i}\hat{w}_{j}+\hat{w}_{i}\hat{n}_{j},$
$e_{ij}^{B}$ $=\hat{m}_{i}\hat{m}_{j}+\hat{n}_{i}\hat{n}_{j},$ $e_{ij}^{L}$ $=\hat{w}_{i}\hat{w}_{j}.$
And the detector tensor $D^{ij}$ can be writen as [60]
$D^{ij}=\frac{1}{2}[\hat{u}^{i}\hat{u}^{j}\mathcal{T}(f,\hat{u}\cdot\hat{w})-\hat{v}^{i}\hat{v}^{j}\mathcal{T}(f,\hat{v}\cdot\hat{w})],$ (8)
with
$\mathcal{T}(f,\hat{a}\cdot\hat{b})=$ $\frac{1}{2}\left\{\sin c\left[\frac{f}{2f_{*}}(1-\hat{a}\cdot\hat{b})\right]\right]$ (9)
×$\times\exp\left[-i\frac{f}{2f_{*}}\left(3+\hat{a}\cdot\hat{b}\right)\right]$
$+\operatorname{sinc}\left[\frac{f}{2f_{*}}\left(1+\hat{a}\cdot\hat{b}\right)\right]$
×$\times\exp\biggl{[}-i\frac{f}{2f_{*}}(1+\hat{a}\cdot\hat{b})\biggr{]}\biggr{\}},$
where $\operatorname{sinc}(x)=\sin x/x$, $f_{*}=c/(2\pi L)$ is the transfer frequency and $L$ is the arm length of the detector. The transfer frequencies for space-based detectors are 19 mHz for LISA, 16 mHz for Taiji, and 275 mHz for TianQin. Due to the arm length, the transfer frequency of ground-based detectors is significantly higher than their sensitivity frequency band.
The angular response functions $F^{A}(\lambda,\beta,\psi)$ or $F^{A}(\alpha,\delta,\psi)$ are obtained by substituting the detector and ecliptic/equatorial coordinates into Eqs. (6)-(9) using Euler rotation conversion. To present the angular response function concisely, we introduce the combined tensor and vector modes:
$F^{T}=\sqrt{\left|F^{+}\right|^{2}+\left|F^{\times}\right|^{2}},$ (10)
$F^{V}=\sqrt{\left|F^{X}\right|^{2}+\left|F^{Y}\right|^{2}}.$
According to Ref. [27], the angular response functions of combined tensor mode $F^{T}$, combined vector mode $F^{V}$, breathing mode $F^{B}$, and longitudinal mode $F^{L}$ are independent of the polarization angle $\psi$. In this section, we study the response of different modes using these four independent modes of $\psi$ and the six basic modes for subsequent research and calculations. On this basis, we calculate the angular response functions for two different frequencies in the LISA detector coordinates to plot Fig. 5.
According to Eq. (9), at the low-frequency limit $f\ll f_{*}$, $\mathcal{T}\to 1$ leads to $F^{B}=-F^{L}$ [27]. Ground-based detectors operate within sensitivity frequency bands below the low-frequency limit, making it impossible to distinguish between these two modes. Moreover, space-based detectors may break through the low-frequency limit and resolve this degeneracy. Figure. 5 shows that at the low-frequency limit, the angular response functions for the breathing mode and longitudinal mode are degenerate at every position. Beyond the transfer frequency, the optimal response position of the breathing mode shifts from $\phi_{d}=45^{\circ},135^{\circ},225^{\circ},315^{\circ}$ to $\phi_{d}=90^{\circ},270^{\circ}$. This shift, which differs from the longitudinal mode’s optimal response position, makes it possible to break the degeneracy of these two modes.
As shown in Fig. 5, the optimal response positions of the combined tensor mode and the combined vector mode change with frequency. The optimal position for the combined tensor mode shifts from a direction perpendicular to the constellation plane to a direction closer to it. For the combined vector mode, the optimal positions at $\phi_{d}=90^{\circ}$ and $\phi_{d}=270^{\circ}$ disappear as frequency increases. In summary, at frequencies beyond the low-frequency limit, there is no degeneracy among the modes, and the optimal response positions generally are not overlapped.
Figure 5 illustrates the response at two specific frequencies. To further examine the relationship between response and frequency, we introduce the averaged angular response function averaged over the source locations [27]:
$R_{A}(f)=\frac{1}{4\pi}\int_{0}^{2\pi}\int_{0}^{\pi}|F^{A}|^{2}\sin\theta_{d}\mathrm{d}\theta_{d}\mathrm{d}\phi_{d},$ (11)
where $A=\{T,V,B,L\}$ denotes four modes independent of $\psi$. We also introduce effective strain noise to measure the detector’s sensitivity to different modes varying with frequency:
$h^{A}_{eff}(f)=\sqrt{\frac{S_{n}(f)}{R_{A}(f)}},$ (12)
where $S_{n}(f)$ is the noise PSD of the detector.
Figure 6: The averaged response functions $R_{A}(f)$ and effective strain noise $h^{A}_{eff}(f)$ of LISA’s four modes vary with frequency. We calculate $R_{A}(f)$ and $h^{A}_{eff}(f)$ for each frequency point through integration. The purple vertical dashed line represents LISA’s transfer frequency $f_{*}=19.1\ \mathrm{mHz}$, and the black dashed line is a fitting approximation of $R_{A}(f)$ (cf. Ref. [27]).
After our calculations, we derive the averaged response functions and effective strain noise for LISA’s four modes, as shown in Fig. 6. At the low-frequency limit, the averaged response function remains constant, with $R_{T}=R_{V}=2(\sin^{2}\gamma)/5$ and $R_{B}=R_{L}=(\sin^{2}\gamma)/15$. As the transfer frequency approaches, $R_{A}$ starts to decrease, showing three distinct damping trends. $R_{T}$ and $R_{V}$ diverge with increasing frequency, similar to $R_{B}$ and $R_{L}$. Additionally, $R_{T}$ and $R_{B}$ show differences at the low-frequency limit. As the frequency increases, the values of those two modes tend to converge. Note that while the average value of these two modes is the same, their responses at individual positions differ, preventing degeneracy between $R_{T}$ and $R_{B}$.
Based on Ref. [27], we analyze the detector’s capability to detect various polarizations using the response function. For ground-based detectors, the sensitive frequency band is below the transfer frequency, preventing the distinction between the breathing mode and the longitudinal mode. Space-based detectors, however, can break this degeneracy at non-low-frequency limits. In Sec. V, we simulate GW signals to evaluate the performance of ground- and space-based detectors in constraining polarization.
IV Methodology
IV.1 Data analysis
Figure 7: Multiband observation of BBHs with ground- and space-based detectors. The vertical axis represents the dimensionless characteristic strain $\sqrt{fS_{n}(f)}$. The BBH curve is derive from Ref. [53] and is intended for illustration purposes only, not for calculation purposes. For luminosity distance selection, we set SBBH at $z=0.01\ (D_{L}=44.6\ \mathrm{Mpc})$ and MBHB at $z=1\ (D_{L}=6.8\ \mathrm{Gpc})$. The cyan line represents the space detection source, while the olive line represents the ground detection source. The three vertical dashed lines represent the transfer frequencies of the three space-based detectors.
In general, the SNR $\rho$ of GW can be defined as
$\rho^{2}=(h|h),$ (13)
where the inner product (|)($\cdot$|$\cdot$)( ⋅ | ⋅ ) generalizes the time-domain correlation product and is conventionally defined as
$(a|b)=4\text{Re}\left[\int_{0}^{\infty}\frac{\tilde{a}^{*}(f)\tilde{b}(f)}{S_{n}(f)}\mathrm{d}f\right],$ (14)
where $\tilde{a}(f)$ and $\tilde{b}(f)$ are the Fourier transforms of $a(t)$ and $b(t)$ respectively. In our study, we consider fourteen GW parameters in total, including ppE parameters, which are
$\bm{\xi}=\{t_{c},m_{1},m_{2},z,\iota,\Phi_{0},\phi_{e},\theta_{e},$ (15)
$\psi,\alpha_{Q},\alpha_{D},\alpha_{V},\alpha_{B},\alpha_{L}\},$
where $t_{c}$ is the coalescence time, $z$ is the redshift of the source, $(\phi_{e},\theta_{e})$ is the sky position, representing ecliptic $(\lambda,\beta)$ or equatorial $(\alpha,\delta)$ coordinates.
For assessing the limitations of the detector on different polarizations and the uncertainty in estimating all parameters, we use the FIM method, defined as
$\Gamma_{ij}=\left(\frac{\partial h}{\partial\xi_{i}}\right|\frac{\partial h}{\partial\xi_{j}}\biggr{)},$ (16)
where $\xi_{i}$ represents the parameter in Eq. (15). For high SNR, the inverse of the FIM, $\Sigma=\Gamma^{-1}$, is the variance-covariance matrix, with the diagonal elements representing variance [61, 62]. Thus, the uncertainty $\Delta\xi_{i}$ of the parameters is given by
$\Delta\xi_{i}=\sqrt{\Sigma_{ii}}.$ (17)
When calculating the FIM in Eq. (16), we use the numerical differentiation approximation from Refs. [44, 63]. Additionally, for the observation network, the total SNR and FIM are obtained by summing the inner products calculated by each detector.
IV.2 BBH source selection
Space- and ground-based detectors have different sensitive frequency bands, resulting in the detection of various GW sources. For BBH sources, ground-based detectors primarily observe SBBHs capturing the complete CBC process of three phases, whereas space-based detectors mainly observe that of MBHBs. Additionally, space-based detectors can detect SBBHs, as many SBBHs inspiral in the low-frequency band before merger, entering the sensitivity band of these detectors. The different BBH sources we selected are presented in Fig. 7.
For all CBC processes, since additional polarizations contribute more significantly in the inspiral phase and Eq. (2) does not apply to the merger and ringdown phases, our waveform model concentrates on the inspiral phase before the binary reaches the innermost stable circular orbit (ISCO). The frequency of ISCO is given by [36]
$f_{\mathrm{ISCO}}=\frac{c^{3}}{6\sqrt{6}\pi GM}.$ (18)
Ground-based detectors typically observe SBBH GWs entering the frequency band a few seconds to minutes before merger. Thus, we select three typical SBBHs with equal mass ratios and total masses of $3\ \mathrm{M_{\odot}}$, $20\ \mathrm{M_{\odot}}$, and $100\ \mathrm{M_{\odot}}$, respectively. We calculate the waveforms for the 10 minutes leading up to ISCO. For MBHB sources observed by space-based detectors, we select MBHBs with three typical masses of $10^{5}\ \mathrm{M_{\odot}}$, $10^{6}\ \mathrm{M_{\odot}}$, and $10^{7}\ \mathrm{M_{\odot}}$, calculating waveforms for the 3 months before reaching ISCO. For multiband observations, we extend two types of SBBHs with masses of $20\ \mathrm{M_{\odot}}$ and $100\ \mathrm{M_{\odot}}$ to the low-frequency band. Those waveforms last for a year before reaching 0.1 Hz. After 0.1 Hz, those SBBHs would take 3 months ($20\ \mathrm{M_{\odot}}$) and 6 days ($100\ \mathrm{M_{\odot}}$) to enter the ground-based detector observation frequency band, providing an opportunity to study multiband observations.
Table 1: Parameter distribution used in calculation [45]. $U[a,b]$ represents a uniform distribution from $a$ to $b$.
Parameter Distribution
Ecliptic Longitude $\lambda$ $U[0,2\pi]$ rad
Ecliptic Latitude $\beta$ $\arcsin(U[-1,1])$ rad
Equatorial Longitude $\alpha$ $U[0,2\pi]$ rad
Equatorial Latitude $\delta$ $\arcsin(U[-1,1])$ rad
Polarization $\psi$ $U[0,2\pi]$ rad
Initial Phase $\phi_{0}$ $U[0,2\pi]$ rad
For the selection of sky positions for BBH sources, we use identical parameters for each source, generating 100 distributions consistent with Table 1. To study the influence of inclination angle, we consider 45 different inclination angles, resulting in 4500 calculations of SNR and FIM for each BBH source, with the results presented in Secs. V.1 and V.2. Finally, for selecting fiducial values for the ppE parameter, we set it to $10^{-4}$ (cf. [36]).
V Constraints on parameters
V.1 Results with inclination
We specifically examine the relationship between the uncertainty of ppE parameters and the inclination angle. According to Eq. (2), we can define the distribution of inclination of the tensor mode $p_{T}(\iota)$ as
$p_{T}(\iota)\propto\sqrt{\left((1+\cos^{2}\iota)/2\right)^{2}+\left(\cos\iota\right)^{2}}.$ (19)
Similarly, we can determine that the distributions of vector mode $p_{V}(\iota)$ and scalar mode $p_{S}(\iota)$ are
$p_{V}(\iota)$ $\propto\sqrt{\left(\cos\iota\right)^{2}+1^{2}},$ (20)
$p_{S}(\iota)$ $\propto\sqrt{\left(\sin\iota\right)^{2}+\left(\sin\iota\right)^{2}}.$
We present the distributions of those three modes in Fig. 8.
Figure 8: The distributions of the inclination for tensor mode $p_{T}(\iota)$, vector mode $p_{V}(\iota)$ and scalar mode $p_{S}(\iota)$. All distributions have been normalized.
Figure 9: Relationship between SNR and parameter uncertainty with inclination angle variation. The vertical axis is represented in logarithmic scale. SNR is calculated based on the GR case, and each line in the graph represents the median value. The results for $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ are identical, so only one is shown. The left panel includes LISA (solid line) and TianQin (dashdot line), while the right panel includes LIGO (solid line) and ET (dashdot line).
From Fig. 8, it is evident that the distribution of scalar mode with inclination is opposite to that of tensor and vector modes, reaching its maximum value at $\iota=\pi/2$ and minimum at $\iota=0,\pi$. From the perspective of inclination angle influence alone, the scalar mode is superior to the other two modes at $\iota=\pi/2$, while tending towards 0 at $\iota=0,\pi$. Tensor and vector modes exhibit the same trend of change, with tensor modes showing a greater amplitude of variation. Furthermore, we calculate the SNR and FIM to obtain the results for the uncertainty of the ppE parameters, as shown in Fig. 9.
As the distribution in Fig. 8, the variation of the scalar mode with inclination in Fig. 9 is opposite to other modes. Furthermore, due to the distribution of scalar mode being zero at $\iota=0,\pi$, there is almost no signal near those two values, resulting in the parameter uncertainty of the scalar mode in Fig. 9 approaching infinity. In addition, considering the low-frequency limit, the $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ of a $10^{5}\ \mathrm{M_{\odot}}$ MBHB with low-SNR are smaller than that of a $10^{7}\ \mathrm{M_{\odot}}$ MBHB with high-SNR. Moreover, due to the superior sensitivity of LISA compared to TianQin and the lower transfer frequency of LISA, the $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ of LISA are significantly better than that of TianQin. For the degeneracy caused by arm length, ground-based detectors cannot distinguish between the breathing mode and the longitudinal mode, so the $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ are much greater than that of space-based detectors and becomes invalid.
Besides the scalar mode, a higher SNR signifies a more robust signal, typically leading to reduced parameter uncertainties. In our calculations, SNR is calculated based on the GR case, where the contribution of tensor modes outweighs that of other modes. Thus, the SNR value is primarily influenced by the tensor modes. In contrast to tensor and scalar modes, the uncertainty of the vector mode is notably higher at $\iota=\pi/2$ in TianQin. This is because TianQin’s fixed orientation restricts the response function’s variability at that specific angle, leading to a pronounced increase in $\Delta\alpha_{V}$ near $\iota=\pi/2$.
The transfer frequency not only directly impacts the degeneracy between breathing and longitudinal modes but also affects other aspects. The transfer function $\mathcal{T}$ is directly related to frequency, providing constraints beyond the low-frequency limit. $\alpha_{Q}$ and $\alpha_{D}$ determine the variation of the orbital angular frequency, and the influence of $\mathcal{T}$ can reduce parameter uncertainty. Therefore, the $\Delta\alpha_{Q}$ of $10^{5}\ \mathrm{M_{\odot}}$ MBHB is smaller than that for a $10^{6}\ \mathrm{M_{\odot}}$ MBHB, and $\Delta\alpha_{D}$ for both MBHBs also varies with the relationship near $\iota=\pi/2$. Because of the low-frequency limit, ground-based detectors are not affected by $\mathcal{T}$, causing the uncertainties $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$ for SBBHs to depend entirely on SNR. This section focuses on presenting and analyzing the relationship between the uncertainty of ppE parameters and the inclination angle for two typical space-based and ground-based detectors.
V.2 Results with ppE parameters
Figure 10: SNR and parameter uncertainty with space- and ground-based detectors. The vertical axis is represented in logarithmic scale. The upper and lower horizontal lines in each box represent the 90% confidence interval. The edges of the box correspond to the upper and lower quartiles, while the lines inside the box represent the median. The uncertainty of sky position is expressed as $\Delta\Omega=2\pi|\sin\theta_{e}|(\Sigma_{\theta_{e}\theta_{e}}\Sigma_{\phi_{e}\phi_{e}}-\Sigma^{2}_{\theta_{e}\phi_{e}})^{1/2}$ [64]. The left panel shows the results for the space-based detectors, where TJ represents Taiji and TQ represents TianQin. In our calculations, there is no difference between the two alternative orbital configurations for TianQin, so they are not labeled separately. The right panel displays the results for the ground-based detectors, but due to the extremely poor positioning capabilities of individual Virgo and KAGRA detectors, their $\Delta\Omega$ values are not shown.
We evaluate the performance of several typical BBHs using different detectors and various parameters. The results for different detectors and their networks are illustrated in Fig. 10.
Regarding space-based detectors, Taiji, which shares a similar configuration with LISA, surpasses LISA across all metrics, demonstrating superior SNR and reduced parameter uncertainty. Moreover, TianQin’s distinct orbital configuration results in inferior performance compared to LISA and Taiji. Additionally, the significant rise in TianQin’s parameter uncertainty at $\iota=\pi/2$, particularly in $\Delta\alpha_{V}$, leads to the upper limit of the box plot for $\Delta\alpha_{V}$ being substantially higher than its median value in the comprehensive results presented in Fig. 10.
When considering the space-based detector network, significant improvements are observed across all metrics compared to individual detectors. Since both LISA and Taiji utilize heliocentric orbits, their network configuration remains stable, maintaining a consistent angle between the detectors. LISA and the three alternative orbital configurations of Taiji exhibit different angles, which introduce variations in the outcomes. A larger angular separation between the detectors correlates with greater coverage of the sky area with high sensitivity, resulting in differences in parameter uncertainty under similar SNR conditions. Overall, the LISA+TJm configuration achieves the most favorable results, with LISA+TJp surpassing LISA+TJc, especially noticeable in parameters like $\Delta\alpha_{V}$ and $\Delta\Omega$. Moreover, it is evident that the LISA+TJ combination outperforms LISA+TQ, and a network of three detectors surpasses two. For more detailed analysis of space detector networks, refer to Ref. [44].
Single ground-based detectors such as Virgo and KAGRA do not perform well in detecting additional polarization parameters and accurately determining the sky position of the source. This limitation arises because a single detector has only one response angle, and a 10-minute GW signal is relatively short compared to the ground-based detector’s observational period (Earth’s rotation period). In the ground-based detector’s coordinate system, a 10-minute SBBH source travels only $2.5^{\circ}$, whereas a 90-day MBHB source moves nearly $90^{\circ}$ in the space-based detector coordinate system. Consequently, the parameter uncertainty range for ground-based detectors is larger. Furthermore, both LIGO and ET have three detectors, but ET detectors are all coplanar, whereas LIGO’s are not. This situation is similar to the superiority of LISA+TJm over LISA+TJc, where the relative positions of LIGO’s three detectors offer greater advantages compared to ET. Therefore, despite the SNR of ET is one order of magnitude higher than LIGO, their $\Delta\Omega$ values are still very similar.
Regarding the ground-based detector network, currently operational LVK cannot fully compensate for the sensitivity differences between second-generation and third-generation detectors. Compared to ET, LVK exhibits a difference of one to two orders of magnitude in the limitations of ppE parameters. Moreover, due to the varied response angles of LVK, $\Delta\Omega$ values are slightly better than those of ET. The LVK+ET combination is the most comprehensive, significantly enhancing network sensitivity. For instance, $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$ can approach the sensitivity level of space-based detector networks, with $\Delta\alpha_{V}$ potentially surpassing them. Due to the arm length, the $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ of the ground-based detector network is four to six orders of magnitude higher than that of the space-based detector network, making it impossible to distinguish between breathing and longitudinal modes.
V.3 Results with redshift
In our previous calculations and analysis, we assume fixed redshift: $z=0.01$ for SBBH and $z=1$ for MBHB. Different redshifts also lead to variations in the GW signal, potentially impacting the final results. From Eq. (2), the strain is inversely proportional to the luminosity distance, represented as $h\propto 1/D_{L}\sim 1/z$. Roughly speaking, as redshift increases, the strain decreases. Regarding the ppE parameter $\xi$, the partial derivative of the strain with respect to it, $\partial h/\partial\xi\propto 1/z$. Consequently, the FIM is $(\partial h/\partial\xi|\partial h/\partial\xi)\propto 1/z^{2}$, and thus the uncertainty of the ppE parameter scales as $\Delta\xi\propto z$. In other words, the uncertainty of the ppE parameter is directly proportional to the redshift:
$\Delta\xi(z)=\mathcal{K}z+\Delta\xi(0),$ (21)
where $\mathcal{K}$ is a proportionality constant. We calculate the results for SBBH and MBHB at different redshifts, as shown in Table 2.
Table 2: Proportional coefficient $\mathcal{K}$ corresponding to different ppE parameters. We use logarithmic $\log(\mathcal{K})$ to represent the results in the table.
Detector $M[\mathrm{M_{\odot}}]$ $\Delta\alpha_{Q}$ $\Delta\alpha_{D}$ $\Delta\alpha_{V}$ $\Delta\alpha_{B}(\alpha_{L})$
$10^{5}$5.01 -8.59 -2.50 -0.82
LISA $10^{6}$4.72 -8.84 -2.42 -0.35
$10^{7}$5.45 -10.13 -2.60 1.62
$10^{5}$5.30 -8.90 -2.79 -1.00
Taiji $10^{6}$4.90 -9.06 -2.52 -0.66
$10^{7}$5.74 -10.42 -2.81 1.36
$10^{5}$4.91 -8.53 -2.25 0.92
TianQin $10^{6}$4.43 -8.68 -1.98 2.32
$10^{7}$5.24 -9.95 -1.93 3.79
$3$1.90 -5.71 0.35 7.93
LIGO $20$2.42 -6.83 -0.18 7.41
$100$2.52 -6.54 -0.15 7.38
$3$1.58 -5.42 2.56 10.09
Virgo $20$2.13 -6.55 2.60 9.59
$100$2.19 -6.21 2.88 9.18
$3$1.44 -5.30 2.91 9.59
KAGRA $20$2.11 -6.57 2.82 8.44
$100$1.89 -5.93 3.26 7.80
$3$2.94 -6.62 -0.85 7.08
ET $20$3.37 -7.78 -1.29 6.62
$100$3.71 -7.81 -1.49 6.30
From Eq. (21), it is clear that the uncertainty of the ppE parameters is proportional to the redshift. Additionally, the rate of change in uncertainty with redshift varies for different ppE parameters. Table 2 illustrates that the influence of redshift on $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$ is negligible, as most $\mathcal{K}$ values are less than $10^{-4}$. That is because these two parameters are directly related to frequency, and the change in GW amplitude due to redshift has a negligible effect on frequency resolution. Moreover, ground-based detectors exhibit relatively larger $\mathcal{K}$ values compared to space-based detectors.
The $\mathcal{K}$ for $\Delta\alpha_{V}$ is significantly larger compared to the first two parameters. Space-based detectors typically have $\mathcal{K}$ values around $10^{-2}$. In contrast, ground-based detectors like Virgo and KAGRA have $\mathcal{K}$ values around $10^{2}$. LIGO and ET perform better, with $\mathcal{K}<1$ in most cases. Those observations are consistent with the analysis in last section, explaining these differences based on detector performance.
Due to the degeneracy in ground-based detectors, $\Delta\alpha_{B}$ ($\Delta\alpha_{L}$) exhibits significantly large $\mathcal{K}$ values, exceeding $10^{6}$. In contrast, space-based detectors, which break this degeneracy, yield noticeably better results. The relationship between the frequency range of MBHBs and the transfer frequency determines the results in Table 2. Smaller mass MBHBs exhibit lower frequencies closer to the transfer frequency, enhancing the resolution of these two modes and leading to smaller $\mathcal{K}$ values. The results are similar for different detectors: Taiji has the smallest transfer frequency, while TianQin has the largest, corresponding to their respective $\mathcal{K}$ values.
V.4 Multiband observation
Space-based detectors can not only observe MBHBs, but also SBBHs. Among tens of thousands of sources, a small number of SBBHs increase in frequency rapidly and merge into the ground-based detector frequency band within a short period, enabling multiband observations by both space- and ground-based detectors [65]. We calculate FIM for two typical SBBHs observed across multiple frequency bands and compared their results with those observed only by ground-based detectors, as shown in Fig. 11.
Figure 11: The uncertainty of ppE parameters in multiband observations. The vertical axis is represented in logarithmic scale. For simplicity, we use points to represent the median, with error bars above and below indicating the 90% confidence interval. In the left panel, the total mass of the SBBH source is $M=20\ \mathrm{M_{\odot}}$, while in the right panel, it is $M=100\ \mathrm{M_{\odot}}$. To save computational resources, only a subset of the sources are calculated, resulting in slight differences from Fig. 10.
Multiband observations provide a significant enhancement to ground-based detectors, particularly in resolving the degeneracy between the breathing and longitudinal modes. The addition of space-based detectors effectively breaks the degeneracy, allowing multiband SBBH observations to limit $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ to levels comparable with space-based detector observations of MBHB. Additionally, multiband observations greatly improve the performance of Virgo and KAGRA in $\Delta\alpha_{V}$, reducing them by nearly three orders of magnitude, approaching the levels achieved by LIGO and ET. Furthermore, the $\Delta\alpha_{V}$ level of the ground-based detector network remains higher than that of the space-based detector network, so the improvement from multiband observations is not as significant.
Regarding $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$, multiband observations have varying improvements for SBBH depending on the mass. The results for SBBH with $M=20\ \mathrm{M_{\odot}}$ do not show significant improvement, while those for $M=100\ \mathrm{M_{\odot}}$ show some improvement. Due to the high sensitivity of ET, multiband observation does not significantly improve ET and LVK+ET, it significantly enhances the performance of only second-generation detectors. Within the space-based detector frequency band, although both types of SBBH have a duration of one year, their frequency variation ranges differ. Specifically, the frequency variation for $M=20\ \mathrm{M_{\odot}}$ SBBHs is 46 mHz, whereas for $M=100\ \mathrm{M_{\odot}}$ it increases to 78 mHz. Greater frequency variations lead to stronger constraints, thereby affecting the observed improvements. Moreover, longer observation times result in larger frequency variations. While four-year observations generally outperform one-year observations, this improvement is not significant. This is because the further away from the merger, the smaller the frequency change over the same time period. For example, in calculations for SBBH with $M=100\ \mathrm{M_{\odot}}$, the frequency change in one year leading up to 0.1 Hz is 78 mHz, but the change over four years is only 87 mHz, indicating a minimal additional change of 9 mHz over three years. Hence, our choice of a one-year observation duration remains reasonable.
The performance of different space-based detectors significantly affects the results of multi band observations, for example, Taiji is superior to LISA. Moreover, the selected cutoff frequency of SBBHs within the space-based detector band is 0.1 Hz, lower than TianQin’s 0.28 Hz transfer frequency. That disparity renders TianQin less effective than LISA and Taiji in observing breathing and longitudinal modes, resulting in a multiband observation improvement for $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ that is two orders of magnitude weaker compared to the other detectors. Conversely, TianQin exhibits significantly enhanced multiband observations for $\Delta\alpha_{Q}$, $\Delta\alpha_{D}$, and $\Delta\alpha_{V}$, because TianQin’s sensitivity within the selected SBBH frequency band surpasses that of the other detectors, leading to superior results. Overall, the addition of space-based detectors compensates for the degeneracy of ground-based detectors in scalar modes and also improves other aspects, enabling the observation results of SBBH to reach the level of MBHB results.
V.5 Multimessenger observation
Apart from multiband observations using space- and ground-based detectors, multimessenger observations, which include the assistance of electromagnetic (EM) observations, can similarly enhance the performance of GW observations. EM observations offer a unique perspective about BBH sources distinct from GW observations. References [66, 67, 68, 69] have demonstrated that mergers of MBHBs can emit EM radiation from accretion disks, while mergers of SBBHs in active galactic nuclei can produce jets, making both scenarios promising targets for EM detectors. Accurate determination of the source’s sky position enables subsequent EM follow-up observations to search for counterparts. The results of EM observations can then serve as valuable priors to reduce uncertainties in GW observations.
The EM effects produced by BBH mergers are expected to be detected by infrared, optical, and X-ray observatories, and the results from these observatories have different impacts on the enhancement of GW. For simplicity, we consider an ideal scenario where EM observations accurately determine specific parameters, ignoring differences between EM detectors. According to Refs. [63, 44], when considering the improvement from EM observations, the corresponding row and column in the FIM are removed to reduce the uncertainties of other parameters in the GW data. That method assesses the performance enhancement of GW observations under ideal multimessenger conditions. We quantify these enhancements when EM observations perfectly determine redshift, sky position, or inclination angle, comparing them against results from GW observations alone. Notably, although this study only focuse on ppE parameters, multimessenger observations also offer varying degrees of enhancement for time, mass, polarization, and other parameters. For details on enhancements in non-ppE parameters, refer to Ref. [44].
For LISA and Taiji, multimessenger observations demonstrate notable enhancements primarily for MBHBs with $M=10^{7}\ \mathrm{M_{\odot}}$. Determining the redshift or inclination angle through multimessenger observations can reduce $\Delta\alpha_{V}$, $\Delta\alpha_{B}$, and $\Delta\alpha_{L}$ by 15% to 24%. Determining the sky position can reduce these parameters by 36% to 49%. Multimessenger observations yield even more significant improvements for TianQin, enhancing detection capabilities across all three typical MBHB masses. When the redshift, inclination angle, or sky position is accurately determined, $\Delta\alpha_{V}$ can be reduced by 40% to 57%. Furthermore, due to arm length, reductions in $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ are more modest, ranging from 9% to 17%.
For ground-based detectors, the impact of multimessenger observations on $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ is considerably smaller compared to multiband observations, as it does not resolve the degeneracy and thus is not considered further. LIGO shows less significant improvement compared to ET, due to its less diverse response angles. Determining the redshift or inclination angle results in a reduction of $\Delta\alpha_{V}$ by 5% to 7%, with negligible changes observed when determining the sky position. In contrast, ET exhibits more substantial enhancements. When any one of the three parameters is determined, $\Delta\alpha_{V}$ can be reduced by 15% to 43%, with larger reductions for higher-mass SBBHs. Moreover, multimessenger observations yield remarkable improvements for Virgo and KAGRA, achieving reductions in $\Delta\alpha_{V}$ by 97% to 99.4%. This improvement is only one order of magnitude lower than that of LIGO and significantly surpasses the outcomes from GW observations alone.
Multimessenger observations, whether for space- or ground-based detectors, do not improve the other two ppE parameters, $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$. That is because these two ppE parameters are directly tied to GW frequency, where GW observation precision is highest. Theoretically, multimessenger observations could enhance constraints on all ppE parameters, but in practice, some improvements are too marginal. Therefore, we only present results where the parameter uncertainty is reduced by more than 5%, with most unreported results being less than 1%. In summary, multimessenger observations can moderately enhance the performance of GW observations, providing stronger constraints on ppE parameters.
VI Conclusion
In this paper, we investigate the expected targets for constraining GW polarization using space- and ground-based detectors within the ppE framework. Specifically, we adopt a model-independent ppE framework that incorporates a GW waveform with all six polarization modes, adhering to current GR test results. For space-based detectors, we consider LISA, Taiji, and TianQin, along with alternative orbital configurations, both individually and in network scenarios. For ground-based detectors, we include the currently operational LVK and the third-generation detector ET, evaluated both individually and in network. Our analysis focuses on detector performance across different polarization modes, emphasizing the response function’s perspective. Furthermore, by simulating three typical masses of MBHB and SBBH, we use the FIM to quantify ppE parameter constraints for both space- and ground-based detectors, presenting the results under different combinations. Furthermore, we explore how multiband and multimessenger observations enhance ppE constraints, offering a comprehensive analysis from diverse angles.
For space-based detectors, Taiji provides the best constraints on GW polarization modes, with LISA performing better than TianQin. Specifically, TianQin’s constraints on scalar modes are significantly weaker than those of LISA and Taiji, while the uncertainties in other ppE parameters among the three detectors differ by less than an order of magnitude. In network scenarios, the combination of LISA+TJm performs slightly better than other combinations, and the network of all three detectors outperforms two-detector networks.
For ground-based detectors, the differences in constraints on frequency-related ppE parameters are minor, but LIGO’s results for vector modes are substantially better than those of Virgo and KAGRA, approaching the level of ET. In network scenarios, the LVK combination does not surpass the results of ET alone, with the LVK+ET network providing the best constraints. Overall, ground-based detectors are unable to distinguish between breathing and longitudinal modes and are significantly less capable of constraining scalar modes compared to space detectors. Other than this, the constraints on ppE parameters from space- and ground-based detectors are comparable.
For multiband observation, it can effectively break the degeneracies of ground-based detectors in breathing and longitudinal modes, addressing the inability of ground detectors to distinguish scalar modes. They also improve the ability of Virgo and KAGRA to constrain vector modes, reaching the level of LIGO, and offer certain improvements for other ppE parameters. Moreover, TianQin brings the greatest improvement, followed by Taiji, and LISA has the least improvement.
For multimessenger observation, it provides the most significant enhancement for TianQin, bringing its results for vector modes close to those of LISA. For scalar modes, LISA and Taiji see greater improvements. ET shows a larger improvement compared to LIGO, and Virgo and KAGRA experience substantial enhancements, with their results for vector modes approaching the level of LIGO. Overall, both multiband and multimessenger observations can enhance GW detection results to a certain extent, leading to better constraints on GW polarization modes.
In future research, we plan to approach the study from several aspects. First, we will use longer-duration GW signals, which can contain more information, and consider the non-light-speed propagation caused by different modes, where vector and scalar modes may arrive before or simultaneously with tensor modes [70]. Second, we aim to delve into the effects brought by Time-Delay Interferometry (TDI) technology, as different TDI combinations may result in subtle differences in the outcomes [71]. Additionally, we will consider various sources, examining the impact of ppE parameters on unequal mass BBH systems. Lastly, in the context of multimessenger observations, we will study the performance of EM detectors under realistic conditions, providing results that are more aligned with actual scenarios to enhance GW observations. In summary, through these in-depth studies, we aim to advance the research on GW polarization modes, evaluate detector performance more comprehensively and in greater detail, and provide more valuable results.
Acknowledgements.
This work was supported by the National Key Research and Development Program of China (Grant No. 2021YFC2203004), the National Natural Science Foundation of China (Grant No. 12347101), and the Natural Science Foundation of Chongqing (Grant No. CSTB2023NSCQ-MSX0103).

### v2

@@
 Abstract
-In general relativity, gravitational waves (GWs) exhibit two tensor modes, while alternative theories predict up to six polarization modes.
-We investigate GW polarization constraints using a model-independent parametrized post-Einsteinian framework with space- and ground-based detectors.
-Evaluating LISA, Taiji, TianQin, LIGO, Virgo, KAGRA, and Einstein Telescope (ET), we analyze their capabilities and network performance.
-Taiji provides the best constraints among space-based detectors, with LISA outperforming TianQin.
-Among ground-based detectors, LIGO excels in vector modes, while ET offers the most comprehensive constraints.
+The theory of general relativity (GR) predicts the existence of gravitational waves (GWs) with two tensor modes, while alternative theories propose up to six polarization modes.
+In this study, we investigate constraints on GW polarization using a model-independent parametrized post-Einsteinian framework and consider both space- and ground-based detectors.
+By evaluating the capabilities and network performance of LISA, Taiji, TianQin, LIGO, Virgo, KAGRA, and the Einstein Telescope (ET), we analyze their respective contributions.
+Among space-based detectors, Taiji provides the most stringent constraints compared with LISA and TianQin.
+Regarding ground-based detectors, LIGO excels in vector modes while ET offers comprehensive constraints across all polarization modes.
 In network scenarios, LISA+TJm performs best, and ET surpasses second-generation detector combinations.
-Multiband observations alleviate scalar mode degeneracies, significantly enhancing ground-based detector performance.
-Combined space- and ground-based observations yield robust constraints on GW polarization, advancing tests of deviations from general relativity.
-Our findings highlight the potential of future GW missions to refine our understanding of gravitational physics through precise polarization measurements.
+Furthermore, multiband observations effectively mitigate scalar mode degeneracies thereby significantly enhancing the performance of ground-based detectors.
+Ultimately, combined space- and ground-based observations provide robust constraints on GW polarizations that advance tests for deviations from GR.
+Our findings underscore the potential of future GW missions in refining our understanding of gravitational physics through precise measurements of polarizations.
 I Introduction
-Over the past century, general relativity (GR) has been rigorously tested through numerous solar system experiments and astrophysical observations, with no definitive evidence suggesting deviations from GR [1, 2, 3].
-Despite passing all experimental tests, there are indications that GR may need to be extended, particularly due to the challenges posed by dark energy and dark matter in cosmology [4, 5].
-Gravitational waves (GWs) offer a new avenue for testing GR, especially in the strong-field regime where GWs are generated by compact objects [6].
+Over the past century, general relativity (GR) has undergone rigorous testing through numerous solar system experiments and astrophysical observations, yielding no definitive evidence suggesting deviations from GR [1, 2, 3].
+Despite successfully passing all experimental tests, there are still indications that GR may require extension due to the challenges from dark energy and dark matter in cosmology [4, 5].
+Gravitational waves (GWs) provide a new way to test GR, especially in the strong-field regime where GWs originate from compact objects [6].
 One method to test GR is by detecting additional polarizations of GWs.
-In GR, GWs only have two tensor modes, whereas some alternative gravity theories predict up to six polarization modes [7].
-For instance, Brans-Dicke theory predicts an additional scalar mode [2, 8], while $f(R)$ theory and Horndeski theory propose two additional scalar modes [9, 10, 11].
+In GR, GWs possess only two tensor modes; however, certain alternative gravity theories predict at most six types of polarizations [7].
+For instance, Brans-Dicke theory predicts an extra scalar mode [2, 8], while $f(R)$ theory and Horndeski theory propose two additional scalar modes [9, 10, 11].
 Einstein-Aether theory predicts five polarization modes [12, 13], and some tensor-vector-scalar theories encompass all six polarization modes [14, 15].
-Therefore, detecting additional polarizations of GWs can help identify deviations from GR, surpass current experimental limits, and reveal the deeper nature of gravity.
-Since LIGO first detected GW [16], nearly a hundred GW events generated by compact binary coalescences (CBCs) have been observed [17, 18, 19].
-Testing GR with these GW data has shown that all observed data are consistent with GR so far [20, 21, 22, 23, 24, 25, 26].
-With the commencement of the fourth observation run (O4), more GW events are expected to be detected, allowing for further testing of GR.
-In general, to detect the polarization of transient GWs, more detectors are needed to obtain responses from different directions, thereby enhancing detection capabilities [27].
-Ground-based detectors in the hertz frequency band, such as LIGO [28], Virgo [29], and KAGRA [30], collaborate to form a detector network to detect the polarization of GWs.
-For continuous GWs emitted by distant CBCs, a detector with directional change capability is sufficient [27].
-Space-based GW detectors in the millihertz frequency band, such as LISA [31], Taiji [32], and TianQin [33], consist of three spacecraft orbiting the Sun/Earth in a triangular formation, making it possible to detect additional GW polarizations due to the detector’s motion in space [34].
-Besides determining whether the detector can theoretically detect additional polarizations of GWs, it is also crucial to ascertain whether the detected GW signal is consistent with the predictions of GR.
-One approach is to use a waveform model with purely phenomenological parameters to represent possible deviations from GR and to apply constraints to these parameters using observed GW data [6].
-Yunes and Pretorius proposed the parameterized post-Einstein (ppE) waveform [35], based on the post-Newtonian (PN) approximation, which is suitable for parameterizing the influence of alternative gravity theories on non-GR polarizations.
-Most modified gravity theories, such as Brans-Dicke theory, massive gravity, and bimetric theory, can be described using the ppE framework [36].
-Although the ppE framework cannot parameterize all possible deviations from GR, it can test for additional polarizations, providing a method to evaluate the performance of detectors in testing GR [37, 38, 39].
-Currently, some work utilizes the ppE framework to constrain additional polarizations for testing GR.
-Narikawa and Tagoshi discussed the potential of advanced ground-based GW detectors, such as LIGO, Virgo, and KAGRA, in detecting general deviations between GWs and the predictions of GR using the ppE framework [40].
+Therefore, detecting additional polarizations of GWs can aid in identifying deviations from GR beyond current experimental limits and unraveling the profound nature of gravity.
+Since the first detection of GW by LIGO in GW150914 [16], nearly a hundred GW events generated by compact binary coalescences (CBCs) have been observed [17, 18, 19].
+The analysis of these GW data has consistently demonstrated that all observed events are in agreement with GR so far [20, 21, 22, 23, 24, 25, 26].
+The start of the fourth observation run (O4), is expected to lead to the detection of more GW events, offering additional opportunities for testing GR.
+It is crucial to use multiple detectors capable of capturing signals from various directions to effectively detect the polarization of transient GWs and enhance detection capabilities.
+[27].
+Ground-based detectors operating in the hertz frequency band, such as LIGO [28], Virgo [29], and KAGRA [30], collaborate to form a network for measuring the polarization of GWs.
+For continuous GWs emitted by distant CBCs, a detector with directional change capability alone would suffice for accurate measurements [27].
+Space-based GW detectors operating in the millihertz frequency band, including LISA [31], Taiji [32], and TianQin [33].
+They consist of three spacecraft forming a triangular formation, orbiting either the Sun or Earth.
+These configurations enable the detection of additional polarizations due to the change in direction [34].
+Theoretically, the capability of a detector to observe additional polarizations of GWs should be assessed, along with verifying the consistency of the detected GW signal with GR predictions.
+One approach involves utilizing a waveform model with purely phenomenological parameters to represent potential deviations from GR and applying constraints to these parameters using observed GW data [6].
+Yunes and Pretorius introduced the parametrized post-Einstein (ppE) waveform [35], derived from the post-Newtonian (PN) approximation, to effectively parametrize the impact of alternative gravity theories on non-GR polarizations.
+Most modified gravity theories, such as Brans-Dicke theory, massive gravity, and bimetric theory, can be characterized within the ppE framework [36].
+The ppE framework cannot parametrize all possible deviations from GR, but it does offer a method for testing additional polarizations [37, 38, 39].
+At present, the ppE framework is being employed to impose constraints on additional polarizations for testing GR.
+Narikawa and Tagoshi explored the potential of advanced ground-based detectors, such as LIGO, Virgo, and KAGRA, in detecting deviations beyond GR using the ppE framework [40].
+Takeda et al.
+estimated the accuracy of polarization determination for CBCs using the Fisher information matrix (FIM) [41, 42].
 Huwyler et al.
 employed LISA to detect massive black hole binaries (MBHBs) and test GR by representing ppE waveforms with phase corrections only [43].
-References [36, 6] discussed the potential of Taiji and TianQin in using MBHBs to test GR and provide constraints on non-GR parameters.
-For detector networks, Nair et al.
-focused on the synergistic effect of the Einstein Telescope (ET) and preDECIGO, demonstrating enhanced sensitivity within a specific band [42].
+The potential of Taiji and TianQin in testing GR and constraining non-GR parameters using MBHBs has been discussed in Refs.
+[36, 6].
+Nair et al.
+investigated the synergistic impact of the Einstein Telescope (ET) and preDECIGO on detector networks, demonstrating improved sensitivity within a specific frequency range [44].
+Xie et al.
+utilized GW signals from nearly ten thousand double white dwarfs observed by the TianQin and LISA networks to detect additional polarization modes [45].
 In Ref.
-[43], Wang and Han used the LISA-Taiji network to test the ppE parameters of deviations from the GR waveform, indicating a significant improvement in the detection of polarization amplitude through joint observations.
+[46], Wang and Han utilized the LISA-Taiji network to investigate the ppE parameters of deviations from the GR waveform, demonstrating a significant enhancement in the detection of polarization amplitude through combined observations.
 Based on our previous work [47, 48, 49], we investigate the constraints from binary black holes (BBHs) for alternative configurations of space- and ground-based detectors on the detection of additional GW polarizations.
-We employ the ppE framework for model-independent testing in GR by numerically calculating time-domain GW signals that contain all polarizations.
-Space-based detectors LISA, Taiji, and TianQin are used to observe massive black hole binaries (MBHBs), while ground-based detectors LIGO, Virgo, KAGRA, and the Einstein Telescope (ET) observe stellar-mass binary black holes (SBBHs), providing constraint results under different networks.
-We also consider the combination of multiband observations.
-The long-term observation of SBBHs by space-based detectors helps break the response degeneracy of ground-based detectors, thereby improving constraints.
-Using the Fisher information matrix (FIM), we present the constraint results of several typical mass BBHs as a function of redshift.
-Additionally, we discuss the potential effects of multimessenger observations on enhancing parameter constraints.
-Through systematic research, we comprehensively analyze the results of additional GW polarizations constrained by space- and ground-based detectors from multiple perspectives.
+The ppE framework is employed for model-independent testing in GR through the numerical calculation of time-domain GW signals that encompass all polarizations.
+The space-based detectors LISA, Taiji, and TianQin are employed for the observation of MBHBs, while the ground-based detectors LIGO, Virgo, KAGRA, and ET observe stellar binary black holes(SBBHs).
+These different networks provide constraint results under distinct observational platforms.
+The combination of multiband observations is also taken into consideration.
+The long-term observation of SBBHs by space-based detectors effectively resolves the response degeneracy in ground-based detector, enhancing their precision.
+The FIM is used to present the constraint results for several typical mass BBHs based on luminosity distance.
+The potential impacts of multimessenger observations on improving parameter constraints are also discussed.
+Through a systematic research approach, we comprehensively analyze the outcomes derived from additional GW polarizations that are constrained by both space- and ground-based detectors, considering multiple perspectives.
 This paper is organized as follows.
@@
 In Sec.
-III, we review the performance parameters and alternative configurations of space- and ground-based detectors, and analyze the response functions corresponding to different modes.
+III, we evaluate the performance and configurations of space and ground detectors, analyzing their response functions for various modes.
 In Sec.
@@
 The GW waveform of an MBHB with a total mass of $10^{5}\ \mathrm{M_{\odot}}$ is observed by LISA over the year before coalescence.
-To clearly visualize differences, we set $\alpha_{D}=0.001$, $\alpha_{Q}=96/5$, and $\alpha_{V}=\alpha_{B}=\alpha_{L}=0.15$, ensuring $\mathcal{A}_{V}=\mathcal{A}_{B}=\mathcal{A}_{L}$.
+To highlight differences, we set $\alpha_{D}=0.001$, $\alpha_{Q}=96/5$, and $\alpha_{V}=\alpha_{B}=\alpha_{L}=0.15$, ensuring $\mathcal{A}_{V}=\mathcal{A}_{B}=\mathcal{A}_{L}$.
 (a) displays the specific time-domain GW signal.
@@
 (d) illustrates the ratio of the amplitude of additional modes to the tensor modes.
-In generalized modified gravity theories, metric perturbations can have up to six independent degrees of freedom, resulting in six different polarization modes: two tensor modes ($+$ and $\times$), two vector modes ($X$ and $Y$) and two scalar modes ($B$ and $L$[47].
+In generalized modified gravity theories, GWs can have at most six types of polarizations: two tensor modes ($+$ and $\times$), two vector modes ($X$ and $Y$) and two scalar modes ($B$ and $L$[50].
 For detectors, the observed GW strain can be described as a linear combination of different GW polarizations, expressed as
@@
 where $A=\{+,\times,X,Y,B,L\}$ represents for the six polarizations, $h_{A}(t)$ is the input signal of GWs, and $F^{A}$ is the angular response function.
-The extended ppE framework is used to construct a model-independent test for GR, including all GW polarization modes [37].
-The amplitude and phase of GWs can be obtained separately from the measurement of perturbations and energy evolution, and there exist simultaneous quadrupole and dipole radiation.
+The extended ppE framework is utilized to construct a model-independent test for GR, encompassing all GW polarization modes [37].
+The amplitude and phase of GWs can be obtained from the metric perturbation and energy evolution, respectively [36].
 The process of CBC includes three phases: inspiral, merger, and ringdown.
-Currently, the observed GW events are transient and do not include the early inspiral phase [19].
-Moreover, GR has passed all current experimental tests, indicating that only tensor modes generated by quadrupole radiation are present in the non-inspiral phase.
-The contribution of dipole radiation in the early inspiral phase can be greater than during the merger phase [36, 48].
-Therefore, it is reasonable to assume that tensor modes dominate in quadrupole radiation, while vector and scalar modes dominate in dipole radiation.
+The current observations of GW events are transient and do not encompass the early inspiral phase [19].
+The successful passing of all current experimental tests in GR provides evidence for the dominance of tensor modes during the merger and ringdown phases.
+The contribution of dipole radiation in the early inspiral phase may exceed that during the merger phase [36, 51].
+The dominance of tensor modes in quadrupole radiation and the dominance of vector and scalar modes in dipole radiation can be reasonably assumed.
 Following Refs.
@@
 where $\alpha_{V,B,L}$ are the dimensionless ppE parameters, $\mathcal{M}=(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$ is the chirp mass, $m_{1}$ and $m_{2}$ are the masses of BBH, $D_{L}$ is the luminosity distance, $\iota$ is the inclination angle, $\Phi=\int{\omega}\mathrm{d}t$ is the orbital phase, $\omega$ is the orbital angular frequency, $\Phi_{0}$ is the initial orbital phase, $G$ and $c$ are the gravitational constant and the speed of light.
-We only consider the dominant modes in different radiations for the amplitude, and the contribution of different radiations to the orbital angular frequency varies.
-Considering the influence of possible non-dominant modes, the overall evolution of the orbital angular frequency can be described by the contributions of dipole radiation and quadrupole radiation [49, 36]:
+The overall evolution of orbital angular frequency can be described by considering the combined effects of dipole and quadrupole radiation [52, 36]
 $\frac{\mathrm{d}\omega}{\mathrm{d}t}=\alpha_{D}\eta^{2/5}\frac{G\mathcal{M}}{c^{3}}\omega^{3}+\alpha_{Q}\left(\frac{G\mathcal{M}}{c^{3}}\right)^{5/3}\omega^{11/3},$ (4)
 where $\alpha_{D}$ and $\alpha_{Q}$ are the ppE parameters that describe the orbital angular frequency contributions of dipole and quadrupole radiation, $\eta=m_{1}m_{2}/M^{2}$ is the symmetric mass ratio, and $M=m_{1}+m_{2}$ is the total mass.
-To describe the model-independent GW waveform, we select five dimensionless ppE parameters to constrain parameters in alternative gravity theories.
+The model-independent GW waveform is described by selecting five dimensionless ppE parameters to constrain the parameters in alternative theories of gravity.
 In the case of GR, $\alpha_{D,V,B,L}=0$ and $\alpha_{Q}=96/5$.
-To ensure our analysis remains independent of specific models, we treat these ppE parameters as independent.
-Some theories exhibit correlations between these parameters, potentially enhancing parameter constraints [36].
-Thus, our results are conservative, and the actual constraints may exceed those presented here.
-From Eq. (4), solving the evolution function $\omega(t)$ of the orbital angular frequency over time analytically is challenging. We determine $t(\omega)$ through integration::
+The independence of our analysis is guaranteed by treating these ppE parameters as mutually independent.
+Theories often show correlations among these parameters, potentially enhancing the constraints on them [36].
+The findings presented here are conservative, and the actual results may be more positive.
+The analytical solution for the time evolution function $\omega(t)$ of the orbital angular frequency, as indicated by Eq. 4), poses a challenging task. We determine $t(\omega)$ through integration,
 $t=t_{0}+\int_{\omega(t_{0})}^{\omega(t)}{\left(\frac{\mathrm{d}\omega}{\mathrm{d}t}\right)^{-1}}\mathrm{d}\omega,$ (5)
@@
 [52, 36].
-Computational methods, such as the bisection method, iteratively solve for the orbital angular frequency corresponding to a given time point.
-By solving point by point, we obtain $\omega(t)$.
+The bisection method and other computational techniques are employed to iteratively determine the orbital angular frequency corresponding to a given time point.
+By solving point by point, we obtain the value of $\omega(t)$.
 Using the above method, we input the calculated $\omega(t)$ into Eqs.
-(1)-(3) to derive the final GW signal.
-In order to better demonstrate the impact of ppE parameters on GW waveforms significantly, we set large parameter values in Fig.
+(1$\sim$(3) to derive the final GW signal.
+In order to better demonstrate the impact of ppE parameters on GW waveforms significantly, we have employed the parameters with large values in Fig.
 1.
 As shown in Fig.
-1(b), there is a notable amplitude difference between the ppE waveform and the GR waveform.
-This discrepancy arises because the GW frequency from dipole radiation equals the orbital frequency of the BBH, whereas quadrupole radiation operates at twice that frequency.
-This effectively superimposes a waveform with half the frequency onto the original GR waveform, resulting in amplitude modulation at twice the wavelength.
-Comparing Figs.
-1(b) and (c), the disparity between the ppE waveform and the GR waveform is significant when the merger is distant but diminishes as the merger approaches, consistent with our assumption in Eq.
-(2) to align with current GR testing outcomes.
-In Fig.
-1(d), the ratio of additional modes to tensor modes is visually depicted, showing a gradual decrease over time, causing the ppE waveform to progressively approach the GR waveform.
-From Eq.
-(3), it can be seen that the amplitude of the tensor modes varies with the power of $2/3$ of the angular frequency, while the additional modes varies with the power of $1/3$.
-The rate of amplitude increase for tensor modes is significantly greater than for additional modes, resulting in the ratio $\mathcal{A}_{V}/\mathcal{A}_{T}\propto\omega^{-1/3}$ declining as angular frequency increases.
+1(b), a noticeable disparity in amplitude is observed between the ppE waveform and the GR waveform.
+The discrepancy arises because the GW frequency resulting from dipole radiation is equivalent to the orbital frequency of the BBH, while quadrupole radiation operates at the twice frequency.
+Consequently, this effectively overlays a waveform with a half of the frequency onto the original GR waveform, leading to amplitude modulation at the twice wavelength.
+The disparity between the ppE waveform and the GR waveform is significant when the merger is far away but diminishes as the merger approaches, which aligns with our assumption in Eq.
+(2) and current GR testing outcomes (see Figs.
+1(b) and (c)).
+The visual representation in Fig.
+1(d) shows a gradual decrease in the ratio of additional modes to tensor modes over time, causing the ppE waveform to converge towards the GR waveform.
+The amplitude of tensor modes varies with an angular frequency raised to the power of $2/3$, while additional modes vary with an angular frequency raised to the power of $1/3$ according to Eq.
+(3).
+The amplitude of tensor modes increases more quickly than that of the additional modes, resulting in a declining ratio $\mathcal{A}_{V}/\mathcal{A}_{T}\propto\omega^{-1/3}$ as angular frequency increases.
 III Detectors and response
@@
 Figure 2: Configuration and location of space- and ground-based detectors (not to scale).
-(a) and (b) depict alternative orbital configurations of LISA, Taiji-p, Taiji-c, Taiji-m, TianQin I, and TianQin II [44].
+(a) and (b) depict alternative orbital configurations of LISA, Taiji-p, Taiji-c, Taiji-m, TianQin I, and TianQin II [53, 54, 55, 56, 47].
 The solid arrow on the constellation plane represents the normal direction of the detector constellation plane.
@@
 This paper uses the obliquity of the ecliptic ($\varepsilon_{2000}=23^{\circ}26^{\prime}21^{\prime\prime}$) based on epoch J2000.0 for the conversion between these two coordinate systems [59].
-In the 2030s, LISA, Taiji, and TianQin are scheduled to launch triangular constellations consisting of three spacecraft.
-LISA and Taiji utilize heliocentric orbits, whereas TianQin employs geocentric orbits.
-Various alternative orbital configurations are available, as detailed in Figs.
+The LISA, Taiji, and TianQin missions are planned to launch triangular constellations comprising three spacecraft in the 2030s.
+LISA and Taiji will adopt heliocentric orbits, while TianQin will employ geocentric orbits.
+Various alternative orbital configurations are available, as illustrated in Figs.
 2 and 3.
-LISA and Taiji, including three configurations, have a leading or trailing angle of $20^{\circ}$ relative to Earth, with a $60^{\circ}$ inclination between the constellation plane and the ecliptic plane [31, 32].
-Furthermore, the normal direction of the TianQin constellation plane remains fixed, and TianQin follows a “three months on+three months off” observation scheme [33].
-Variations in the constellation plane determine how GW sources appear in the detector coordinate system, thereby influencing the response to different polarizations.
-Additionally, detector sensitivity is directly influenced by the arm length, with LISA, Taiji, and TianQin having arm lengths of $2.5\times 10^{6}$ km, $3\times 10^{6}$ km, and $\sqrt{3}\times 10^{5}$ km, respectively.
-Concerning the noise in space-based detectors, we account for acceleration noise, displacement noise, and foreground noise.
-Acceleration noise and displacement noise contribute to the primary power spectral density (PSD), as detailed in Refs.
+The LISA and Taiji missions exhibit a leading or trailing angle of $20^{\circ}$ with respect to Earth, while maintaining a $60^{\circ}$ inclination between the constellation plane and the ecliptic plane [31, 32].
+The TianQin constellation plane remains fixed in its normal direction and follows an observation scheme characterized by alternating periods of “three months on+three months off” [33].
+The orientation of the constellation plane affects how GW sources appear in the detector coordinate system and influences their response to different polarizations.
+The sensitivity of the detectors is directly affected by the length of their arms.
+LISA, Taiji, and TianQin have arm lengths of $2.5\times 10^{6}$ km, $3\times 10^{6}$ km, and $\sqrt{3}\times 10^{5}$ km respectively.
+The noise in space-based detectors includes acceleration, displacement, and foreground noise.
+The primary power spectral density (PSD) includes contributions from acceleration and displacement noises, as described in Refs.
 [60, 61, 62].
-Foreground noise originates from galactic binaries within the Milky Way galaxy, creating a peak in the frequency band of approximately 0.3-3 mHz, as demonstrated in our previous work [44].
+The foreground noise comes from galactic binaries in our Milky Way galaxy, creating a peak frequency band of 0.3-3 mHz [47].
 III.2 Ground-based detectors
-In contrast to the planned space-based detectors, the ground-based detectors are currently operational, primarily comprising the LVK collaboration, which includes LIGO, Virgo and KAGRA.
-These are all L-shaped second-generation detectors with arm lengths of 4 km (LIGO), 3 km (Virgo), and 3 km (KAGRA), respectively [28, 29, 30].
-In this paper, we consider LIGO to include three detectors: LIGO Hanford (H1) and LIGO Livingston (L1), which are operational, and the planned LIGO India (I1).
-Furthermore, we consider the third-generation detector ET, which features a triangular shape composed of three 10 km arms [56].
-As the Earth rotates, ground-based detectors scan different sky regions, as illustrated in Fig.
+Unlike the planned space-based detectors, the operational ground-based detectors primarily include the LVK Collaboration comprising LIGO (4 km), Virgo (3 km), and KAGRA (3 km) [28, 29, 30].
+These detectors are all second-generation with an L-shaped configuration.
+In this paper, we consider LIGO as comprising three operational detectors: LIGO Hanford (H1), LIGO Livingston (L1), and the planned LIGO India (I1).
+The third-generation detector ET is characterized by a triangular shape with three arms, each measuring 10 km [63].
+The rotation of the Earth causes ground-based detectors to scan different regions of the sky, as depicted in Fig.
 3.
-Due to the obliquity of the ecliptic, the sky regions scanned by space- and ground-based detectors is not parallel.
-Ground-based detectors, unlike their space-based counterparts, are affected by various types of noise, including quantum noise, seismic noise, gravity-gradient noise, thermal noise, and others.
-For our study, we utilize the design performance specifications of these detectors, and the corresponding PSD can be referenced in LIGO Document T1500293 and Refs.
+Due to the obliquity of the ecliptic, the sky regions scanned by space- and ground-based detectors are not parallel.
+The performance of ground-based detectors is affected by various types of noise, including quantum noise, seismic noise, gravity-gradient noise, thermal noise, and others.
+In our study, we utilize the design specifications for these detectors and refer to the corresponding PSD in LIGO Document T1500293 and Refs.
 [64, 65].
@@
 Figure 5: Angular response functions for the combined tensor mode $F^{T}$, combined vector mode $F^{V}$, breathing mode $F^{B}$, and longitudinal mode $F^{L}$ at LISA detector coordinates $(\phi_{d},\theta_{d})$.
-The left panel shows the response functions for GWs at 1 mHz, where the breathing and longitudinal modes are almost identical.
-The right panel displays the response at twice the transfer frequency, $2f_{*}=38\ \mathrm{mHz}$, where the degeneracy between the breathing and longitudinal modes is broken at higher frequencies.
-A dual-arm Michelson interferometer detects GWs by measuring the relative change in the length of its two arms.
-We describe the detector’s response to GWs using the method outlined in Refs.
+The left panel shows the response functions for GWs at 1 mHz, where the breathing and longitudinal modes exhibit almost indistinguishable characteristics.
+The right panel displays the response at twice the transfer frequency, $2f_{*}=38\ \mathrm{mHz}$, where a distinction between the breathing and longitudinal modes emerges as frequencies increase.
+The detection of GWs by a dual-arm Michelson interferometer is achieved by measuring the relative change in length between its two arms.
+We describe the detector’s response using the method outlined in Refs.
 [48, 27].
-As shown in Fig.
-4, the detector coordinates are constructed using orthogonal unit vectors $\hat{x},\hat{y},\hat{z}$}, and the GW coordinates are constructed using $\hat{p},\hat{q},\hat{w}$}.
-Here, $\hat{w}$ represents the propagation direction of the GW, and the unit vector $\hat{\Omega}=-\hat{w}$ represents the position of the GW source.
-In a triangular detector, the angle $\gamma$ between the two arms is $60^{\circ}$, while in an L-shaped detector, it is $90^{\circ}$.
+The detector coordinates, as illustrated in Fig.
+4, are constructed using orthogonal unit vectors $\hat{x},\hat{y},\hat{z}$}.
+Similarly, the GW coordinates are constructed using $\hat{p},\hat{q},\hat{w}$}.
+Here, $\hat{w}$ represents the propagation direction of the GW, while the unit vector $\hat{\Omega}=-\hat{w}$ denotes the position of the GW source.
+The angle $\gamma$ between the two arms is $60^{\circ}$ in a triangular detector and $90^{\circ}$ in an L-shaped detector.
 For GWs, an additional rotational degree of freedom can be fixed by specifying the polarization angle $\psi$, ultimately using $\hat{m},\hat{n},\hat{w}$} to describe the GW.
@@
 $e_{ij}^{X}$ $=\hat{m}_{i}\hat{w}_{j}+\hat{w}_{i}\hat{m}_{j},$ $e_{ij}^{Y}$ $=\hat{n}_{i}\hat{w}_{j}+\hat{w}_{i}\hat{n}_{j},$
-$e_{ij}^{B}$ $=\hat{m}_{i}\hat{m}_{j}+\hat{n}_{i}\hat{n}_{j},$ $e_{ij}^{L}$ $=\hat{w}_{i}\hat{w}_{j}.$
-And the detector tensor $D^{ij}$ can be writen as [60]
+$e_{ij}^{B}$ $=\hat{m}_{i}\hat{m}_{j}+\hat{n}_{i}\hat{n}_{j},$ $e_{ij}^{L}$ $=\hat{w}_{i}\hat{w}_{j},$
+and the detector tensor $D^{ij}$ can be written as [67]
 $D^{ij}=\frac{1}{2}[\hat{u}^{i}\hat{u}^{j}\mathcal{T}(f,\hat{u}\cdot\hat{w})-\hat{v}^{i}\hat{v}^{j}\mathcal{T}(f,\hat{v}\cdot\hat{w})],$ (8)
 with
-$\mathcal{T}(f,\hat{a}\cdot\hat{b})=$ $\frac{1}{2}\left\{\sin c\left[\frac{f}{2f_{*}}(1-\hat{a}\cdot\hat{b})\right]\right]$ (9)
+$\mathcal{T}(f,\hat{a}\cdot\hat{b})=$ $\frac{1}{2}\biggr{\{}\operatorname{sinc}\left[\frac{f}{2f_{*}}(1-\hat{a}\cdot\hat{b})\right]$ (9)
 ×$\times\exp\left[-i\frac{f}{2f_{*}}\left(3+\hat{a}\cdot\hat{b}\right)\right]$
@@
 The transfer frequencies for space-based detectors are 19 mHz for LISA, 16 mHz for Taiji, and 275 mHz for TianQin.
-Due to the arm length, the transfer frequency of ground-based detectors is significantly higher than their sensitivity frequency band.
-The angular response functions $F^{A}(\lambda,\beta,\psi)$ or $F^{A}(\alpha,\delta,\psi)$ are obtained by substituting the detector and ecliptic/equatorial coordinates into Eqs.
-(6)-(9) using Euler rotation conversion.
+Due to the arm length, the transfer frequency of ground-based detectors exceeds their sensitivity frequency band significantly.
+The angular response functions $F^{A}(\lambda,\beta,\psi)$ or $F^{A}(\alpha,\delta,\psi)$ are derived by substituting the detector and ecliptic/equatorial coordinates into Eqs.
+(6$\sim$(9) using Euler rotation conversion.
 To present the angular response function concisely, we introduce the combined tensor and vector modes:
@@
 $F^{V}=\sqrt{\left|F^{X}\right|^{2}+\left|F^{Y}\right|^{2}}.$
-According to Ref.
-[27], the angular response functions of combined tensor mode $F^{T}$, combined vector mode $F^{V}$, breathing mode $F^{B}$, and longitudinal mode $F^{L}$ are independent of the polarization angle $\psi$.
-In this section, we study the response of different modes using these four independent modes of $\psi$ and the six basic modes for subsequent research and calculations.
-On this basis, we calculate the angular response functions for two different frequencies in the LISA detector coordinates to plot Fig.
+The angular response functions of the combined tensor mode $F^{T}$, combined vector mode $F^{V}$, breathing mode $F^{B}$, and longitudinal mode $F^{L}$ are independent of polarization angle $\psi$ [27].
+The response of different modes is studied using these four independent modes of $\psi$ and the six basic modes for subsequent research and calculations.
+On this basis, we calculate the angular response functions for two different frequencies in the LISA detector coordinates to generate Fig.
 5.
 According to Eq.
-(9), at the low-frequency limit $f\ll f_{*}$, $\mathcal{T}\to 1$ leads to $F^{B}=-F^{L}$ [27].
-Ground-based detectors operate within sensitivity frequency bands below the low-frequency limit, making it impossible to distinguish between these two modes.
-Moreover, space-based detectors may break through the low-frequency limit and resolve this degeneracy.
-Figure.
-5 shows that at the low-frequency limit, the angular response functions for the breathing mode and longitudinal mode are degenerate at every position.
+(9), in the low-frequency regime $f\ll f_{*}$, the limit $\mathcal{T}\to 1$ implies $F^{B}=-F^{L}$ [27].
+The sensitivity frequency bands of ground-based detectors are below this low-frequency limit, making it impossible to distinguish between these two modes.
+The potential of space-based detectors to surpass the low-frequency limit and resolve this degeneracy is significant.
+The angular response functions for the breathing mode and longitudinal mode are degenerate at every position in Fig.
+5, indicating their degeneracy at the low-frequency limit.
 Beyond the transfer frequency, the optimal response position of the breathing mode shifts from $\phi_{d}=45^{\circ},135^{\circ},225^{\circ},315^{\circ}$ to $\phi_{d}=90^{\circ},270^{\circ}$.
-This shift, which differs from the longitudinal mode’s optimal response position, makes it possible to break the degeneracy of these two modes.
+The shift, which is distinct from the optimal response position of the longitudinal mode, allows for the differentiation between these two modes.
 As shown in Fig.
@@
 In summary, at frequencies beyond the low-frequency limit, there is no degeneracy among the modes, and the optimal response positions generally are not overlapped.
-Figure 5 illustrates the response at two specific frequencies.
-To further examine the relationship between response and frequency, we introduce the averaged angular response function averaged over the source locations [27]:
+The response at two specific frequencies is illustrated in Fig.
+5.
+In order to further investigate the relationship between response and frequency, we introduce the angular response function averaged over the source locations [27],
 $R_{A}(f)=\frac{1}{4\pi}\int_{0}^{2\pi}\int_{0}^{\pi}|F^{A}|^{2}\sin\theta_{d}\mathrm{d}\theta_{d}\mathrm{d}\phi_{d},$ (11)
-where $A=\{T,V,B,L\}$ denotes four modes independent of $\psi$. We also introduce effective strain noise to measure the detector’s sensitivity to different modes varying with frequency:
+where $A=\{T,V,B,L\}$ denotes four modes independent of $\psi$. The detector’s sensitivity to different frequency modes can be measured by introducing effective strain noise,
 $h^{A}_{eff}(f)=\sqrt{\frac{S_{n}(f)}{R_{A}(f)}},$ (12)
@@
 We calculate $R_{A}(f)$ and $h^{A}_{eff}(f)$ for each frequency point through integration.
-The purple vertical dashed line represents LISA’s transfer frequency $f_{*}=19.1\ \mathrm{mHz}$, and the black dashed line is a fitting approximation of $R_{A}(f)$ (cf.
+The purple vertical-dashed line represents LISA’s transfer frequency $f_{*}=19.1\ \mathrm{mHz}$, and the black dashed line is a fitting approximation of $R_{A}(f)$ (cf.
 Ref.
 [27]).
-After our calculations, we derive the averaged response functions and effective strain noise for LISA’s four modes, as shown in Fig.
+Through the calculations, we obtain the averaged response functions and effective strain noise for LISA’s four modes, as shown in Fig.
 6.
 At the low-frequency limit, the averaged response function remains constant, with $R_{T}=R_{V}=2(\sin^{2}\gamma)/5$ and $R_{B}=R_{L}=(\sin^{2}\gamma)/15$.
-As the transfer frequency approaches, $R_{A}$ starts to decrease, showing three distinct damping trends.
-$R_{T}$ and $R_{V}$ diverge with increasing frequency, similar to $R_{B}$ and $R_{L}$.
-Additionally, $R_{T}$ and $R_{B}$ show differences at the low-frequency limit.
-As the frequency increases, the values of those two modes tend to converge.
-Note that while the average value of these two modes is the same, their responses at individual positions differ, preventing degeneracy between $R_{T}$ and $R_{B}$.
+The decrease in $R_{A}$ becomes evident as the transfer frequency increases, exhibiting three distinct damping trends.
+$R_{T}$ and $R_{V}$ diverge as the frequency rises, similar to the behavior of $R_{B}$ and $R_{L}$.
+The differences between $R_{T}$ and $R_{B}$ become apparent at low frequencies, but as the frequency increases, these two modes tend to converge.
+The average values of these two modes are identical, but their responses at individual positions differ, preventing degeneracy between $R_{T}$ and $R_{B}$.
 Based on Ref.
 [27], we analyze the detector’s capability to detect various polarizations using the response function.
-For ground-based detectors, the sensitive frequency band is below the transfer frequency, preventing the distinction between the breathing mode and the longitudinal mode.
-Space-based detectors, however, can break this degeneracy at non-low-frequency limits.
+The sensitive frequency band of ground-based detectors is below the transfer frequency, making it difficult to distinguish between the breathing mode and longitudinal mode.
+The degeneracy at nonlow-frequency limits can be resolved by space-based detectors.
 In Sec.
-V, we simulate GW signals to evaluate the performance of ground- and space-based detectors in constraining polarization.
+V, we simulate GW signals to assess the capability of ground- and space-based detectors in polarization constraint.
 IV Methodology
@@
 [60] and is intended for illustration purposes only, not for calculation purposes.
-For luminosity distance selection, we set SBBH at $z=0.01\ (D_{L}=44.6\ \mathrm{Mpc})$ and MBHB at $z=1\ (D_{L}=6.8\ \mathrm{Gpc})$.
+For luminosity distance selection, we set SBBH at $D_{L}=44.6$ Mpc ($z=0.01$ in $\mathrm{\Lambda CDM}$ cosmological model) and MBHB at $D_{L}=6.79$ Gpc ($z=1$).
 The cyan line represents the space detection source, while the olive line represents the ground detection source.
@@
 $(a|b)=4\text{Re}\left[\int_{0}^{\infty}\frac{\tilde{a}^{*}(f)\tilde{b}(f)}{S_{n}(f)}\mathrm{d}f\right],$ (14)
-where $\tilde{a}(f)$ and $\tilde{b}(f)$ are the Fourier transforms of $a(t)$ and $b(t)$ respectively. In our study, we consider fourteen GW parameters in total, including ppE parameters, which are
-$\bm{\xi}=\{t_{c},m_{1},m_{2},z,\iota,\Phi_{0},\phi_{e},\theta_{e},$ (15)
+where $\tilde{a}(f)$ and $\tilde{b}(f)$ are the Fourier transforms of $a(t)$ and $b(t)$, respectively.
+The total number of GW parameters considered in our study is fourteen, including ppE parameters, which are
+$\bm{\xi}=\{t_{c},m_{1},m_{2},D_{L},\iota,\Phi_{0},\phi_{e},\theta_{e},$ (15)
 $\psi,\alpha_{Q},\alpha_{D},\alpha_{V},\alpha_{B},\alpha_{L}\},$
-where $t_{c}$ is the coalescence time, $z$ is the redshift of the source, $(\phi_{e},\theta_{e})$ is the sky position, representing ecliptic $(\lambda,\beta)$ or equatorial $(\alpha,\delta)$ coordinates.
-For assessing the limitations of the detector on different polarizations and the uncertainty in estimating all parameters, we use the FIM method, defined as
+where $t_{c}$ is the coalescence time, $D_{L}$ is the luminosity distance of the source, $(\phi_{e},\theta_{e})$ is the sky position, representing ecliptic $(\lambda,\beta)$ or equatorial $(\alpha,\delta)$ coordinates.
+To evaluate the detectors’ performance on different polarizations, we use the FIM:
 $\Gamma_{ij}=\left(\frac{\partial h}{\partial\xi_{i}}\right|\frac{\partial h}{\partial\xi_{j}}\biggr{)},$ (16)
@@
 $\Delta\xi_{i}=\sqrt{\Sigma_{ii}}.$ (17)
-When calculating the FIM in Eq.
-(16), we use the numerical differentiation approximation from Refs.
+When calculating the FIM, we use the numerical differentiation approximation as described in Refs.
 [47, 70].
-Additionally, for the observation network, the total SNR and FIM are obtained by summing the inner products calculated by each detector.
+The total SNR and FIM for the observation network are obtained by summing the inner products calculated by each detector.
 IV.2 BBH source selection
-Space- and ground-based detectors have different sensitive frequency bands, resulting in the detection of various GW sources.
-For BBH sources, ground-based detectors primarily observe SBBHs capturing the complete CBC process of three phases, whereas space-based detectors mainly observe that of MBHBs.
-Additionally, space-based detectors can detect SBBHs, as many SBBHs inspiral in the low-frequency band before merger, entering the sensitivity band of these detectors.
-The different BBH sources we selected are presented in Fig.
+The distinct sensitive frequency bands of space- and ground-based detectors facilitate the detection of various GW sources.
+Ground-based detectors can observe SBBH’s whole CBC in three phases, while space detectors do the same for MBHB.
+Additionally, space detectors can also observe the inspiraling SBBH in the low-frequency band.
+The selected BBH sources are shown in Fig.
 7.
-For all CBC processes, since additional polarizations contribute more significantly in the inspiral phase and Eq.
-(2) does not apply to the merger and ringdown phases, our waveform model concentrates on the inspiral phase before the binary reaches the innermost stable circular orbit (ISCO).
+The proposed waveform model focuses on the inspiral phase before the binary reaches the innermost stable circular orbit (ISCO).
+Additional polarizations plays a more significant role in this phase than in merger and ringdown phases.
 The frequency of ISCO is given by [36]
 $f_{\mathrm{ISCO}}=\frac{c^{3}}{6\sqrt{6}\pi GM}.$ (18)
-Ground-based detectors typically observe SBBH GWs entering the frequency band a few seconds to minutes before merger.
-Thus, we select three typical SBBHs with equal mass ratios and total masses of $3\ \mathrm{M_{\odot}}$, $20\ \mathrm{M_{\odot}}$, and $100\ \mathrm{M_{\odot}}$, respectively.
-We calculate the waveforms for the 10 minutes leading up to ISCO.
-For MBHB sources observed by space-based detectors, we select MBHBs with three typical masses of $10^{5}\ \mathrm{M_{\odot}}$, $10^{6}\ \mathrm{M_{\odot}}$, and $10^{7}\ \mathrm{M_{\odot}}$, calculating waveforms for the 3 months before reaching ISCO.
-For multiband observations, we extend two types of SBBHs with masses of $20\ \mathrm{M_{\odot}}$ and $100\ \mathrm{M_{\odot}}$ to the low-frequency band.
-Those waveforms last for a year before reaching 0.1 Hz.
-After 0.1 Hz, those SBBHs would take 3 months ($20\ \mathrm{M_{\odot}}$) and 6 days ($100\ \mathrm{M_{\odot}}$) to enter the ground-based detector observation frequency band, providing an opportunity to study multiband observations.
+Ground-based detectors typically observe the entry of SBBH GWs into the frequency band a few seconds to minutes prior to their merger.
+Therefore, we have selected three typical SBBHs with equal mass ratios and total masses of $3\ \mathrm{M_{\odot}}$, $20\ \mathrm{M_{\odot}}$, and $100\ \mathrm{M_{\odot}}$, respectively.
+The waveforms for the 10 minutes preceding ISCO are computed.
+For MBHB sources observed by space-based detectors, we select MBHBs with three representative masses: $10^{5}\ \mathrm{M_{\odot}}$, $10^{6}\ \mathrm{M_{\odot}}$, and $10^{7}\ \mathrm{M_{\odot}}$.
+The waveforms are calculated for the three months before ISCO.
+For multiband observations, the SBBHs with masses of $20\ \mathrm{M_{\odot}}$ and $100\ \mathrm{M_{\odot}}$ are extended to the low-frequency band, lasting for a year before reaching 0.1 Hz.
+The SBBHs beyond this threshold would take approximately three months ($20\ \mathrm{M_{\odot}}$) and six days ($100\ \mathrm{M_{\odot}}$) to enter the frequency band of ground-based detectors, providing an opportunity for conducting multiband observations.
 Table 1: Parameter distribution used in calculation [48]. $U[a,b]$ represents a uniform distribution from $a$ to $b$.
@@
 Initial phase $\phi_{0}$ $U[0,2\pi]$ rad
-For the selection of sky positions for BBH sources, we use identical parameters for each source, generating 100 distributions consistent with Table 1.
-To study the influence of inclination angle, we consider 45 different inclination angles, resulting in 4500 calculations of SNR and FIM for each BBH source, with the results presented in Secs.
+We generate 100 BBH sources.
+The sky positions for BBH sources are selected with identical parameters, following the distribution in Table 1.
+In order to investigate the impact of inclination angle, we consider 45 angles, resulting in a total of 4500 calculations of SNR and FIM for each BBH source.
+The corresponding results are presented in Secs.
 V.1 and V.2.
-Finally, for selecting fiducial values for the ppE parameter, we set it to $10^{-4}$ (cf.
+Finally, we set the fiducial value for the ppE parameter to $10^{-4}$ (cf.
 [36]).
@@
 V.1 Results with inclination
-We specifically examine the relationship between the uncertainty of ppE parameters and the inclination angle.
-According to Eq.
-(2), we can define the distribution of inclination of the tensor mode $p_{T}(\iota)$ as
+The relationship between ppE parameter uncertainty and the inclination angle is specifically examined. By referring to Eq. (2), we can define the distribution of inclination for the tensor mode as
 $p_{T}(\iota)\propto\sqrt{\left((1+\cos^{2}\iota)/2\right)^{2}+\left(\cos\iota\right)^{2}}.$ (19)
-Similarly, we can determine that the distributions of vector mode $p_{V}(\iota)$ and scalar mode $p_{S}(\iota)$ are
+The distributions of vector mode $p_{V}(\iota)$ and scalar mode $p_{S}(\iota)$ can be determined similarly,
 $p_{V}(\iota)$ $\propto\sqrt{\left(\cos\iota\right)^{2}+1^{2}},$ (20)
 $p_{S}(\iota)$ $\propto\sqrt{\left(\sin\iota\right)^{2}+\left(\sin\iota\right)^{2}}.$
-We present the distributions of those three modes in Fig. 8.
+The distributions of these three modes are presented in Fig. 8.
 Figure 8: The distributions of the inclination for tensor mode $p_{T}(\iota)$, vector mode $p_{V}(\iota)$ and scalar mode $p_{S}(\iota)$. All distributions have been normalized.
@@
 The vertical axis is represented in logarithmic scale.
-SNR is calculated based on the GR case, and each line in the graph represents the median value.
-The results for $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ are identical, so only one is shown.
+The SNR is calculated based on the GR case, and each line in the graph represents the median value.
+The results for $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ are identical, therefore only one is shown.
 The left panel includes LISA (solid line) and TianQin (dashdot line), while the right panel includes LIGO (solid line) and ET (dashdot line).
-From Fig.
-8, it is evident that the distribution of scalar mode with inclination is opposite to that of tensor and vector modes, reaching its maximum value at $\iota=\pi/2$ and minimum at $\iota=0,\pi$.
-From the perspective of inclination angle influence alone, the scalar mode is superior to the other two modes at $\iota=\pi/2$, while tending towards 0 at $\iota=0,\pi$.
-Tensor and vector modes exhibit the same trend of change, with tensor modes showing a greater amplitude of variation.
-Furthermore, we calculate the SNR and FIM to obtain the results for the uncertainty of the ppE parameters, as shown in Fig.
+The distribution of the scalar mode with inclination angle, as shown in Fig.
+8, exhibits an opposite trend to that of tensor and vector modes.
+It reaches its maximum value at $\iota=\pi/2$ and minimum values at $\iota=0,\pi$.
+The scalar mode outperforms the other two modes at $\iota=\pi/2$, while approaching zero at $\iota=0,\pi$.
+The patterns of variation in both tensor and vector modes are similar, although tensor modes show a greater amplitude of change.
+The SNR and FIM have been computed to determine the uncertainty of ppE parameters, as shown in Fig.
 9.
-As the distribution in Fig.
-8, the variation of the scalar mode with inclination in Fig.
-9 is opposite to other modes.
-Furthermore, due to the distribution of scalar mode being zero at $\iota=0,\pi$, there is almost no signal near those two values, resulting in the parameter uncertainty of the scalar mode in Fig.
-9 approaching infinity.
-In addition, considering the low-frequency limit, the $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ of a $10^{5}\ \mathrm{M_{\odot}}$ MBHB with low-SNR are smaller than that of a $10^{7}\ \mathrm{M_{\odot}}$ MBHB with high-SNR.
-Moreover, due to the superior sensitivity of LISA compared to TianQin and the lower transfer frequency of LISA, the $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ of LISA are significantly better than that of TianQin.
-For the degeneracy caused by arm length, ground-based detectors cannot distinguish between the breathing mode and the longitudinal mode, so the $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ are much greater than that of space-based detectors and becomes invalid.
-Besides the scalar mode, a higher SNR signifies a more robust signal, typically leading to reduced parameter uncertainties.
-In our calculations, SNR is calculated based on the GR case, where the contribution of tensor modes outweighs that of other modes.
-Thus, the SNR value is primarily influenced by the tensor modes.
-In contrast to tensor and scalar modes, the uncertainty of the vector mode is notably higher at $\iota=\pi/2$ in TianQin.
-This is because TianQin’s fixed orientation restricts the response function’s variability at that specific angle, leading to a pronounced increase in $\Delta\alpha_{V}$ near $\iota=\pi/2$.
+The scalar mode in Fig.
+9 shows an opposite trend compared with other modes, as depicted in the distribution shown in Fig.
+8.
+The scalar mode vanishes at $\iota=0,\pi$, resulting in a near absence of signal around these values.
+As a result, the parameter uncertainty for the scalar mode approaches infinity in Fig.
+9.
+Additionally, considering the low-frequency limit, a $10^{5}\ \mathrm{M_{\odot}}$ MBHB with low-SNR demonstrates smaller $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ values compared with a $10^{7}\ \mathrm{M_{\odot}}$ MBHB with high-SNR.
+Owing to superior sensitivity and the lower transfer frequency, LISA exhibits significantly better $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ values than TianQin does.
+Ground-based detectors are unable to distinguish between the breathing mode and longitudinal mode; hence their corresponding $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ values are much larger than those of space-based detectors and become invalid.
+Besides the scalar mode, a higher SNR indicates a more robust signal with reduced parameter uncertainties.
+In our calculations, SNR is computed based on the GR case.
+Therefore, the SNR value is primarily influenced by tensor modes.
+In contrast to tensor and scalar modes, the uncertainty of the vector mode is significantly higher at $\iota=\pi/2$ for TianQin.
+The fixed orientation of TianQin limits variations in the response function at that specific angle, resulting in a significant increase in $\Delta\alpha_{V}$ near $\iota=\pi/2$.
 The transfer frequency not only directly impacts the degeneracy between breathing and longitudinal modes but also affects other aspects.
@@
 Therefore, the $\Delta\alpha_{Q}$ of $10^{5}\ \mathrm{M_{\odot}}$ MBHB is smaller than that for a $10^{6}\ \mathrm{M_{\odot}}$ MBHB, and $\Delta\alpha_{D}$ for both MBHBs also varies with the relationship near $\iota=\pi/2$.
-Because of the low-frequency limit, ground-based detectors are not affected by $\mathcal{T}$, causing the uncertainties $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$ for SBBHs to depend entirely on SNR.
-This section focuses on presenting and analyzing the relationship between the uncertainty of ppE parameters and the inclination angle for two typical space-based and ground-based detectors.
+Because of the low-frequency limit, ground-based detectors are not affected by $\mathcal{T}$, causing the uncertainties $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$ for SBBHs that depend solely on SNR.
+The aim of this section is to analyze the correlation between ppE parameter uncertainty and inclination angle for two detectors, one space-based and one ground-based.
 V.2 Results with ppE parameters
@@
 The right panel displays the results for the ground-based detectors, but due to the extremely poor positioning capabilities of individual Virgo and KAGRA detectors, their $\Delta\Omega$ values are not shown.
-We evaluate the performance of several typical BBHs using different detectors and various parameters. The results for different detectors and their networks are illustrated in Fig. 10.
-Regarding space-based detectors, Taiji, which shares a similar configuration with LISA, surpasses LISA across all metrics, demonstrating superior SNR and reduced parameter uncertainty.
-Moreover, TianQin’s distinct orbital configuration results in inferior performance compared to LISA and Taiji.
-Additionally, the significant rise in TianQin’s parameter uncertainty at $\iota=\pi/2$, particularly in $\Delta\alpha_{V}$, leads to the upper limit of the box plot for $\Delta\alpha_{V}$ being substantially higher than its median value in the comprehensive results presented in Fig.
+The performance of several typical BBHs is assessed using different detectors and various parameters. The results for different detectors and their networks are depicted in Fig. 10.
+Regarding space-based detectors, Taiji, which shares a similar configuration with LISA, surpasses LISA across all aspects, demonstrating superior SNR and reduced parameter uncertainty.
+TianQin’s unique orbital configuration results in inferior performance compared with LISA and Taiji.
+Additionally, the significant increase in TianQin’s parameter uncertainty at $\iota=\pi/2$, particularly in $\Delta\alpha_{V}$, causes the upper limit of the box plot for $\Delta\alpha_{V}$ to be substantially higher than its median value as shown in the comprehensive results presented in Fig.
 10.
-When considering the space-based detector network, significant improvements are observed across all metrics compared to individual detectors.
-Since both LISA and Taiji utilize heliocentric orbits, their network configuration remains stable, maintaining a consistent angle between the detectors.
-LISA and the three alternative orbital configurations of Taiji exhibit different angles, which introduce variations in the outcomes.
-A larger angular separation between the detectors correlates with greater coverage of the sky area with high sensitivity, resulting in differences in parameter uncertainty under similar SNR conditions.
-Overall, the LISA+TJm configuration achieves the most favorable results, with LISA+TJp surpassing LISA+TJc, especially noticeable in parameters like $\Delta\alpha_{V}$ and $\Delta\Omega$.
-Moreover, it is evident that the LISA+TJ combination outperforms LISA+TQ, and a network of three detectors surpasses two.
-For more detailed analysis of space detector networks, refer to Ref.
+When considering the space-based detector network, significant improvements are observed across all metrics compared with individual detectors.
+Due to the heliocentric orbits for both LISA and Taiji, their network configuration remains stable, ensuring a consistent angle between the detectors.
+The three alternative orbital configurations of Taiji, along with LISA, exhibit distinct angles that introduce outcome variations.
+A larger angular separation between the detectors increases sky area coverage with high sensitivity, resulting in differences in parameter uncertainty under similar SNR conditions.
+The LISA+TJm configuration demonstrates superior performance overall, with LISA+TJp outperforming LISA+TJc, particularly in parameters such as $\Delta\alpha_{V}$ and $\Delta\Omega$.
+Furthermore, it is evident that LISA+TJ+TQ surpasses LISA+TJ, highlighting the advantage of a network comprising three detectors over two.
+For a more comprehensive analysis of space detector networks, please refer to Ref.
 [47].
-Single ground-based detectors such as Virgo and KAGRA do not perform well in detecting additional polarization parameters and accurately determining the sky position of the source.
-This limitation arises because a single detector has only one response angle, and a 10-minute GW signal is relatively short compared to the ground-based detector’s observational period (Earth’s rotation period).
+The ability of ground-based detectors like Virgo and KAGRA to detect additional polarization and determine the sky position is limited.
+This limitation stems from the fact that a single detector only has only one response angle, and a 10-minute GW signal is very short compared with the observational period (Earth’s rotation period).
 In the ground-based detector’s coordinate system, a 10-minute SBBH source travels only $2.5^{\circ}$, whereas a 90-day MBHB source moves nearly $90^{\circ}$ in the space-based detector coordinate system.
 Consequently, the parameter uncertainty range for ground-based detectors is larger.
-Furthermore, both LIGO and ET have three detectors, but ET detectors are all coplanar, whereas LIGO’s are not.
-This situation is similar to the superiority of LISA+TJm over LISA+TJc, where the relative positions of LIGO’s three detectors offer greater advantages compared to ET.
-Therefore, despite the SNR of ET is one order of magnitude higher than LIGO, their $\Delta\Omega$ values are still very similar.
-Regarding the ground-based detector network, currently operational LVK cannot fully compensate for the sensitivity differences between second-generation and third-generation detectors.
-Compared to ET, LVK exhibits a difference of one to two orders of magnitude in the limitations of ppE parameters.
-Moreover, due to the varied response angles of LVK, $\Delta\Omega$ values are slightly better than those of ET.
-The LVK+ET combination is the most comprehensive, significantly enhancing network sensitivity.
-For instance, $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$ can approach the sensitivity level of space-based detector networks, with $\Delta\alpha_{V}$ potentially surpassing them.
-Due to the arm length, the $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ of the ground-based detector network is four to six orders of magnitude higher than that of the space-based detector network, making it impossible to distinguish between breathing and longitudinal modes.
-V.3 Results with redshift
-In our previous calculations and analysis, we assume fixed redshift: $z=0.01$ for SBBH and $z=1$ for MBHB.
-Different redshifts also lead to variations in the GW signal, potentially impacting the final results.
+Furthermore, both LIGO and ET have three detectors, but ET’s detectors are all coplanar while LIGO’s are not.
+The relative positions of LIGO’s three detectors offer greater advantages compared with ET.
+Despite having an SNR one order of magnitude higher than LIGO, ET still has very similar $\Delta\Omega$ values.
+Regarding the ground-based detector network, the currently operational LVK is unable to compensate for the sensitivity differences between second-generation and third-generation detectors.
+The limitations of ppE parameters in LVK differ by one to two orders of magnitude compared with ET.
+Furthermore, due to LVK’s varied response angles, $\Delta\Omega$ values are slightly superior to those of ET.
+The combination of LVK+ET provides the most comprehensive enhancement in network sensitivity.
+For instance, $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$ can approach the sensitivity level achieved by space-based detector networks, with potential surpassing in $\Delta\alpha_{V}$.
+Due to its arm length, the ground-based detector network’s $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ are four to six orders of magnitude higher than those observed in space-based detector networks, making it impossible to distinguish between breathing and longitudinal modes.
+V.3 Results with luminosity distance
+The distances we assumed in our previous calculations and analysis were fixed: $D_{L}=44.6$ Mpc for SBBH and $D_{L}=6.79$ Gpc for MBHB.
+The GW signal can vary due to different distances, potentially affecting the final results.
 From Eq.
-(2), the strain is inversely proportional to the luminosity distance, represented as $h\propto 1/D_{L}\sim 1/z$.
-Roughly speaking, as redshift increases, the strain decreases.
-Regarding the ppE parameter $\xi$, the partial derivative of the strain with respect to it, $\partial h/\partial\xi\propto 1/z$.
-Consequently, the FIM is $(\partial h/\partial\xi|\partial h/\partial\xi)\propto 1/z^{2}$, and thus the uncertainty of the ppE parameter scales as $\Delta\xi\propto z$.
-In other words, the uncertainty of the ppE parameter is directly proportional to the redshift:
-$\Delta\xi(z)=\mathcal{K}z+\Delta\xi(0),$ (21)
-where $\mathcal{K}$ is a proportionality constant. We calculate the results for SBBH and MBHB at different redshifts, as shown in Table 2.
+(2), the strain is inversely proportional to the luminosity distance, represented as $h\propto 1/D_{L}$.
+Regarding the ppE parameter $\xi$, the partial derivative of the strain with respect to it, $\partial h/\partial\xi\propto 1/D_{L}$.
+Consequently, the FIM is $(\partial h/\partial\xi|\partial h/\partial\xi)\propto 1/D_{L}^{2}$, and thus the uncertainty of the ppE parameter scales as $\Delta\xi\propto D_{L}$.
+In other words, the uncertainty of the ppE parameter is directly proportional to the distance,
+$\Delta\xi(D_{L})=\mathcal{K}D_{L}+\Delta\xi(0),$ (21)
+where $\mathcal{K}$ is a proportionality constant. The results for SBBH and MBHB are calculated at various distances, as presented in Table 2.
 Table 2: Proportional coefficient $\mathcal{K}$ corresponding to different ppE parameters. We use logarithmic $\log(\mathcal{K})$ to represent the results in the table.
 Detector $M[\mathrm{M_{\odot}}]$ $\Delta\alpha_{Q}$ $\Delta\alpha_{D}$ $\Delta\alpha_{V}$ $\Delta\alpha_{B}(\alpha_{L})$
-$10^{5}$5.01 -8.59 -2.50 -0.82
-LISA $10^{6}$4.72 -8.84 -2.42 -0.35
-$10^{7}$5.45 -10.13 -2.60 1.62
-$10^{5}$5.30 -8.90 -2.79 -1.00
-Taiji $10^{6}$4.90 -9.06 -2.52 -0.66
-$10^{7}$5.74 -10.42 -2.81 1.36
-$10^{5}$4.91 -8.53 -2.25 0.92
-TianQin $10^{6}$4.43 -8.68 -1.98 2.32
-$10^{7}$5.24 -9.95 -1.93 3.79
-$3$1.90 -5.71 0.35 7.93
-LIGO $20$2.42 -6.83 -0.18 7.41
-$100$2.52 -6.54 -0.15 7.38
-$3$1.58 -5.42 2.56 10.09
-Virgo $20$2.13 -6.55 2.60 9.59
-$100$2.19 -6.21 2.88 9.18
-$3$1.44 -5.30 2.91 9.59
-KAGRA $20$2.11 -6.57 2.82 8.44
-$100$1.89 -5.93 3.26 7.80
-$3$2.94 -6.62 -0.85 7.08
-ET $20$3.37 -7.78 -1.29 6.62
-$100$3.71 -7.81 -1.49 6.30
-From Eq.
-(21), it is clear that the uncertainty of the ppE parameters is proportional to the redshift.
-Additionally, the rate of change in uncertainty with redshift varies for different ppE parameters.
-Table 2 illustrates that the influence of redshift on $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$ is negligible, as most $\mathcal{K}$ values are less than $10^{-4}$.
-That is because these two parameters are directly related to frequency, and the change in GW amplitude due to redshift has a negligible effect on frequency resolution.
-Moreover, ground-based detectors exhibit relatively larger $\mathcal{K}$ values compared to space-based detectors.
-The $\mathcal{K}$ for $\Delta\alpha_{V}$ is significantly larger compared to the first two parameters.
-Space-based detectors typically have $\mathcal{K}$ values around $10^{-2}$.
-In contrast, ground-based detectors like Virgo and KAGRA have $\mathcal{K}$ values around $10^{2}$.
-LIGO and ET perform better, with $\mathcal{K}<1$ in most cases.
-Those observations are consistent with the analysis in last section, explaining these differences based on detector performance.
-Due to the degeneracy in ground-based detectors, $\Delta\alpha_{B}$ ($\Delta\alpha_{L}$) exhibits significantly large $\mathcal{K}$ values, exceeding $10^{6}$.
-In contrast, space-based detectors, which break this degeneracy, yield noticeably better results.
-The relationship between the frequency range of MBHBs and the transfer frequency determines the results in Table 2.
-Smaller mass MBHBs exhibit lower frequencies closer to the transfer frequency, enhancing the resolution of these two modes and leading to smaller $\mathcal{K}$ values.
-The results are similar for different detectors: Taiji has the smallest transfer frequency, while TianQin has the largest, corresponding to their respective $\mathcal{K}$ values.
+$10^{5}$9.05 -12.63 -6.55 -4.86
+LISA $10^{6}$8.76 -12.89 -6.46 -4.39
+$10^{7}$9.50 -14.17 -6.64 -2.42
+$10^{5}$9.35 -12.94 -6.83 -5.05
+TaiJi $10^{6}$8.94 -13.11 -6.57 -4.71
+$10^{7}$9.78 -14.46 -6.85 -2.68
+$10^{5}$8.95 -12.58 -6.30 -3.12
+TianQin $10^{6}$8.47 -12.73 -6.03 -1.72
+$10^{7}$9.28 -13.99 -5.97 -0.25
+$3$5.74 -9.54 -3.49 4.10
+LIGO $20$6.25 -10.67 -4.02 3.57
+$100$6.35 -10.38 -3.98 3.54
+$3$5.41 -9.26 -1.27 6.25
+Virgo $20$5.96 -10.38 -1.24 5.75
+$100$6.03 -10.04 -0.96 5.35
+$3$5.28 -9.14 -0.92 5.76
+KAGRA $20$5.94 -10.40 -1.01 4.61
+$100$5.73 -9.77 -0.58 3.97
+$3$6.77 -10.45 -4.68 3.24
+ET $20$7.20 -11.62 -5.13 2.79
+$100$7.54 -11.64 -5.32 2.46
+Equation (21) reveals that the uncertainty of the ppE parameters is directly proportional to the distance.
+The rate of change in uncertainty with distance varies for different ppE parameters.
+Table 2 illustrates that the influence of distance on $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$ is negligible, since most $\mathcal{K}$ values are less than $10^{-5}$.
+This can be attributed to the fact that these two parameters are closely linked to frequency, and any alteration in GW amplitude caused by distance has an negligible effect on frequency resolution.
+Moreover, ground-based detectors exhibit relatively larger $\mathcal{K}$ values compared with space-based detectors.
+The $\mathcal{K}$ for $\Delta\alpha_{V}$ exhibits a significantly larger magnitude than the first two parameters.
+The $\mathcal{K}$ values of space-based detectors are typically around $10^{-6}$, while Virgo and KAGRA have $\mathcal{K}$ values around $10^{-1}$.
+LIGO and ET outperform Virgo and KAGRA, in most cases with $\mathcal{K}<10^{-3}$.
+These observations are consistent the analysis presented in the previous section, which elucidates these differences stem from detector performance.
+Due to the degeneracy in ground-based detectors, $\Delta\alpha_{B}$ ($\Delta\alpha_{L}$) exhibits significantly large values of $\mathcal{K}$, exceeding $10^{2}$.
+In contrast, space-based detectors effectively resolve this degeneracy and improves outcomes.
+The results in Table 2 depend on the frequency range of MBHBs and the transfer frequency.
+The frequencies of smaller mass MBHBs are lower and closer to the transfer frequency, enhancing the resolution of these two modes and resulting in smaller $\mathcal{K}$ values.
+The results are consistent across different detectors: Taiji exhibits the lowest transfer frequency, whereas TianQin demonstrates the highest, aligning with their respective $\mathcal{K}$ values.
 V.4 Multiband observation
-Space-based detectors can not only observe MBHBs, but also SBBHs.
-Among tens of thousands of sources, a small number of SBBHs increase in frequency rapidly and merge into the ground-based detector frequency band within a short period, enabling multiband observations by both space- and ground-based detectors [65].
-We calculate FIM for two typical SBBHs observed across multiple frequency bands and compared their results with those observed only by ground-based detectors, as shown in Fig.
+Space-based detectors can observe both MBHBs and SBBHs.
+Among numerous sources, only a few SBBHs rapidly increase in frequency and merge into the frequency band of ground-based detectors, enabling multiband observations by both space- and ground-based detectors [72].
+The FIM is calculated for two types of SBBHs observed across multiple frequency bands and compared with results solely obtained from ground-based detectors, as shown in Fig.
 11.
@@
 10.
-Multiband observations provide a significant enhancement to ground-based detectors, particularly in resolving the degeneracy between the breathing and longitudinal modes.
-The addition of space-based detectors effectively breaks the degeneracy, allowing multiband SBBH observations to limit $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ to levels comparable with space-based detector observations of MBHB.
-Additionally, multiband observations greatly improve the performance of Virgo and KAGRA in $\Delta\alpha_{V}$, reducing them by nearly three orders of magnitude, approaching the levels achieved by LIGO and ET.
-Furthermore, the $\Delta\alpha_{V}$ level of the ground-based detector network remains higher than that of the space-based detector network, so the improvement from multiband observations is not as significant.
-Regarding $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$, multiband observations have varying improvements for SBBH depending on the mass.
-The results for SBBH with $M=20\ \mathrm{M_{\odot}}$ do not show significant improvement, while those for $M=100\ \mathrm{M_{\odot}}$ show some improvement.
-Due to the high sensitivity of ET, multiband observation does not significantly improve ET and LVK+ET, it significantly enhances the performance of only second-generation detectors.
-Within the space-based detector frequency band, although both types of SBBH have a duration of one year, their frequency variation ranges differ.
-Specifically, the frequency variation for $M=20\ \mathrm{M_{\odot}}$ SBBHs is 46 mHz, whereas for $M=100\ \mathrm{M_{\odot}}$ it increases to 78 mHz.
-Greater frequency variations lead to stronger constraints, thereby affecting the observed improvements.
-Moreover, longer observation times result in larger frequency variations.
+Multiband observations greatly enhances the capabilities of ground-based detectors, particularly in effectively distinguishing between breathing and longitudinal modes.
+The addition of space-based detectors resolves the degeneracy, enabling multiband observations of SBBH to constrain $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ at levels comparable with those achieved by observing MBHB with space-based detectors.
+Additionally, the multiband observations greatly enhance the performance of Virgo and KAGRA in $\Delta\alpha_{V}$, reducing it by nearly three orders of magnitude.
+This brings them closer to the levels achieved by LIGO and ET.
+The $\Delta\alpha_{V}$ level of the ground-based detector network remains higher than that of the space-based detector network, even with multiband observations.
+Therefore, while there is improvement from these observations, its significance may not be as pronounced.
+The improvements for SBBH due to multiband observations vary with the mass, with respect to $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$.
+For SBBH with $M=20\ \mathrm{M_{\odot}}$, there is no significant improvement observed, while for a mass of $100\ \mathrm{M_{\odot}}$, some enhancement is demonstrated.
+The sensitivity of ET is so high that multiband observation does not significantly improve ET and LVK+ET, but it notably enhances the performance of only second-generation detectors.
+The frequency ranges of both types of SBBH differ within the space-based detector band, despite their one-year duration.
+The frequency variation is 46 mHz for SBBHs with $M=20\ \mathrm{M_{\odot}}$ and 78 mHz for those with $M=100\ \mathrm{M_{\odot}}$.
+The stronger constraints are caused by greater frequency variations, which in turn affect the observed improvements.
+Additionally, larger frequency variations occur with longer observation times.
 While four-year observations generally outperform one-year observations, this improvement is not significant.
 This is because the further away from the merger, the smaller the frequency change over the same time period.
-For example, in calculations for SBBH with $M=100\ \mathrm{M_{\odot}}$, the frequency change in one year leading up to 0.1 Hz is 78 mHz, but the change over four years is only 87 mHz, indicating a minimal additional change of 9 mHz over three years.
-Hence, our choice of a one-year observation duration remains reasonable.
-The performance of different space-based detectors significantly affects the results of multi band observations, for example, Taiji is superior to LISA.
-Moreover, the selected cutoff frequency of SBBHs within the space-based detector band is 0.1 Hz, lower than TianQin’s 0.28 Hz transfer frequency.
-That disparity renders TianQin less effective than LISA and Taiji in observing breathing and longitudinal modes, resulting in a multiband observation improvement for $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ that is two orders of magnitude weaker compared to the other detectors.
-Conversely, TianQin exhibits significantly enhanced multiband observations for $\Delta\alpha_{Q}$, $\Delta\alpha_{D}$, and $\Delta\alpha_{V}$, because TianQin’s sensitivity within the selected SBBH frequency band surpasses that of the other detectors, leading to superior results.
-Overall, the addition of space-based detectors compensates for the degeneracy of ground-based detectors in scalar modes and also improves other aspects, enabling the observation results of SBBH to reach the level of MBHB results.
+The SBBH with $M=100\ \mathrm{M_{\odot}}$ experiences a frequency shift of 78 mHz within one year before reaching 0.1 Hz.
+The change over four years was only 87 mHz, with the additional change within three years being just 9 mHz.
+Hence, our choice of a one-year observation duration remains justifiable.
+The performance of various space-based detectors significantly impacts the outcomes of multi-band observations.
+Moreover, the cutoff frequency of SBBHs within the space-based detector band is chosen to be 0.1 Hz, which is lower than TianQin’s transfer frequency of 0.28 Hz.
+That disparity renders TianQin less effective than LISA and Taiji in observing breathing and longitudinal modes, resulting in a much weaker multiband observation improvement for $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$.
+Conversely, TianQin exhibits significantly enhanced multiband observations for $\Delta\alpha_{Q}$, $\Delta\alpha_{D}$, and $\Delta\alpha_{V}$ due to its superior sensitivity within the selected SBBH frequency band compared with other detectors.
+The inclusion of space-based detectors compensates for the degeneracy observed in ground-based detectors regarding scalar modes and enhances other aspects.
+The advancement allows SBBH observation results to reach a level comparable to MBHB results.
 V.5 Multimessenger observation
-Apart from multiband observations using space- and ground-based detectors, multimessenger observations, which include the assistance of electromagnetic (EM) observations, can similarly enhance the performance of GW observations.
-EM observations offer a unique perspective about BBH sources distinct from GW observations.
-References [66, 67, 68, 69] have demonstrated that mergers of MBHBs can emit EM radiation from accretion disks, while mergers of SBBHs in active galactic nuclei can produce jets, making both scenarios promising targets for EM detectors.
-Accurate determination of the source’s sky position enables subsequent EM follow-up observations to search for counterparts.
-The results of EM observations can then serve as valuable priors to reduce uncertainties in GW observations.
-The EM effects produced by BBH mergers are expected to be detected by infrared, optical, and X-ray observatories, and the results from these observatories have different impacts on the enhancement of GW.
-For simplicity, we consider an ideal scenario where EM observations accurately determine specific parameters, ignoring differences between EM detectors.
+Apart from using space- and ground-based detectors for multiband observations, integrating multimessenger observations, including electromagnetic (EM) observations, can also enhance GW detections.
+EM observations provide a distinctive viewpoint on BBH sources that is separate from GW observations.
+The studies conducted by Refs.
+[73, 74] have demonstrated that mergers of MBHBs can emit EM radiation from accretion disks.
+Additionally, the mergers of SBBHs in active galactic nuclei have the potential to generate powerful EM jets [73, 74, 75, 76].
+These findings highlight the significance of both scenarios as promising targets for EM detectors.
+Accurate determination of the source’s sky position facilitates subsequent EM follow-up observations to identify counterparts.
+The findings from EM observations can then serve as valuable prior information to minimize uncertainties in GW observations.
+The EM effects generated by BBH mergers can be detected by infrared, optical, and x-ray observatories, with diverse implications for advancing GW research.
+For simplicity, we consider an ideal scenario where EM observations accurately determine specific parameters, disregarding discrepancies among EM detectors.
 According to Refs.
-[63, 44], when considering the improvement from EM observations, the corresponding row and column in the FIM are removed to reduce the uncertainties of other parameters in the GW data.
-That method assesses the performance enhancement of GW observations under ideal multimessenger conditions.
-We quantify these enhancements when EM observations perfectly determine redshift, sky position, or inclination angle, comparing them against results from GW observations alone.
-Notably, although this study only focuse on ppE parameters, multimessenger observations also offer varying degrees of enhancement for time, mass, polarization, and other parameters.
-For details on enhancements in non-ppE parameters, refer to Ref.
+[70, 47], the FIM removes the corresponding row and column when incorporating EM observations in order to mitigate uncertainties of other parameters in the GW data.
+This approach evaluates the performance enhancement of GW observations under ideal multimessenger conditions.
+The enhancements are quantified when EM observations accurately determine the distance, sky position, or inclination angle, and these results are compared with GW observations alone.
+Notably, although this study only focuse on ppE parameters, multimessenger observations also enhance time, mass, polarization, and other parameters.
+For details on enhancements in non-ppE parameters, please refer to Ref.
 [47].
-For LISA and Taiji, multimessenger observations demonstrate notable enhancements primarily for MBHBs with $M=10^{7}\ \mathrm{M_{\odot}}$.
-Determining the redshift or inclination angle through multimessenger observations can reduce $\Delta\alpha_{V}$, $\Delta\alpha_{B}$, and $\Delta\alpha_{L}$ by 15% to 24%.
-Determining the sky position can reduce these parameters by 36% to 49%.
-Multimessenger observations yield even more significant improvements for TianQin, enhancing detection capabilities across all three typical MBHB masses.
-When the redshift, inclination angle, or sky position is accurately determined, $\Delta\alpha_{V}$ can be reduced by 40% to 57%.
-Furthermore, due to arm length, reductions in $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ are more modest, ranging from 9% to 17%.
-For ground-based detectors, the impact of multimessenger observations on $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ is considerably smaller compared to multiband observations, as it does not resolve the degeneracy and thus is not considered further.
-LIGO shows less significant improvement compared to ET, due to its less diverse response angles.
-Determining the redshift or inclination angle results in a reduction of $\Delta\alpha_{V}$ by 5% to 7%, with negligible changes observed when determining the sky position.
-In contrast, ET exhibits more substantial enhancements.
-When any one of the three parameters is determined, $\Delta\alpha_{V}$ can be reduced by 15% to 43%, with larger reductions for higher-mass SBBHs.
-Moreover, multimessenger observations yield remarkable improvements for Virgo and KAGRA, achieving reductions in $\Delta\alpha_{V}$ by 97% to 99.4%.
-This improvement is only one order of magnitude lower than that of LIGO and significantly surpasses the outcomes from GW observations alone.
+The multimessenger observations for LISA and Taiji demonstrate significant enhancements primarily for MBHBs with $M=10^{7}\ \mathrm{M_{\odot}}$.
+Utilizing multimessenger observations to determine the distance or inclination angle can result in a reduction of $\Delta\alpha_{V}$, $\Delta\alpha_{B}$, and $\Delta\alpha_{L}$ by 15% to 24%.
+The determination of the sky position can lead to a reduction in these parameters ranging from 36% to 49%.
+The multimessenger observations significantly enhance TianQin’s detection capabilities for all three typical MBHB masses.
+When accurately determining the distance, inclination angle, or sky position, $\Delta\alpha_{V}$ can be reduced by 40% to 57%.
+Furthermore, reductions in $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ are less, ranging from 9% to 17%.
+The impact of multimessenger observations on $\Delta\alpha_{B}$ and $\Delta\alpha_{L}$ is significantly diminished for ground-based detectors compared with multiband observations, as it fails to resolve the degeneracy and thus is not further considered.
+The improvement shown by LIGO is relatively less significant than that of ET, mainly because of its varied response angles.
+The determination of the distance or inclination angle leads to a reduction in $\Delta\alpha_{V}$ by 5% to 7%, while negligible changes are observed when determining the sky position.
+The enhancements exhibited by ET are more significant.
+When any one of the three parameters is determined, $\Delta\alpha_{V}$ can be reduced by 15% to 43%, with greater reductions observed for higher-mass SBBHs.
+Moreover, multimessenger observations significantly enhance Virgo and KAGRA, reducing $\Delta\alpha_{V}$ by 97% to 99.4%.
+The improvement achieved is only one order of magnitude lower than that observed in LIGO, demonstrating a significant advancement beyond GW observations alone.
 Multimessenger observations, whether for space- or ground-based detectors, do not improve the other two ppE parameters, $\Delta\alpha_{Q}$ and $\Delta\alpha_{D}$.
-That is because these two ppE parameters are directly tied to GW frequency, where GW observation precision is highest.
-Theoretically, multimessenger observations could enhance constraints on all ppE parameters, but in practice, some improvements are too marginal.
-Therefore, we only present results where the parameter uncertainty is reduced by more than 5%, with most unreported results being less than 1%.
-In summary, multimessenger observations can moderately enhance the performance of GW observations, providing stronger constraints on ppE parameters.
+The reason is that these two ppE parameters directly link to the highest frequency of GWs, where precision in observations is at its peak.
+Theoretically, multimessenger observations could enhance constraints on all ppE parameters, but in practice, some improvements only offer marginal benefits.
+The results presented in this study are limited to cases where the parameter uncertainty is reduced by more than 5%, with most unreported results reducing less than 1%.
+In conclusion, multimessenger observations can greatly improve the effectiveness of GW observations, resulting in stronger constraints on ppE parameters.
 VI Conclusion
@@
 Specifically, we adopt a model-independent ppE framework that incorporates a GW waveform with all six polarization modes, adhering to current GR test results.
-For space-based detectors, we consider LISA, Taiji, and TianQin, along with alternative orbital configurations, both individually and in network scenarios.
-For ground-based detectors, we include the currently operational LVK and the third-generation detector ET, evaluated both individually and in network.
-Our analysis focuses on detector performance across different polarization modes, emphasizing the response function’s perspective.
-Furthermore, by simulating three typical masses of MBHB and SBBH, we use the FIM to quantify ppE parameter constraints for both space- and ground-based detectors, presenting the results under different combinations.
-Furthermore, we explore how multiband and multimessenger observations enhance ppE constraints, offering a comprehensive analysis from diverse angles.
-For space-based detectors, Taiji provides the best constraints on GW polarization modes, with LISA performing better than TianQin.
-Specifically, TianQin’s constraints on scalar modes are significantly weaker than those of LISA and Taiji, while the uncertainties in other ppE parameters among the three detectors differ by less than an order of magnitude.
-In network scenarios, the combination of LISA+TJm performs slightly better than other combinations, and the network of all three detectors outperforms two-detector networks.
-For ground-based detectors, the differences in constraints on frequency-related ppE parameters are minor, but LIGO’s results for vector modes are substantially better than those of Virgo and KAGRA, approaching the level of ET.
-In network scenarios, the LVK combination does not surpass the results of ET alone, with the LVK+ET network providing the best constraints.
-Overall, ground-based detectors are unable to distinguish between breathing and longitudinal modes and are significantly less capable of constraining scalar modes compared to space detectors.
-Other than this, the constraints on ppE parameters from space- and ground-based detectors are comparable.
-For multiband observation, it can effectively break the degeneracies of ground-based detectors in breathing and longitudinal modes, addressing the inability of ground detectors to distinguish scalar modes.
-They also improve the ability of Virgo and KAGRA to constrain vector modes, reaching the level of LIGO, and offer certain improvements for other ppE parameters.
-Moreover, TianQin brings the greatest improvement, followed by Taiji, and LISA has the least improvement.
+The space-based detectors considered in this study include LISA, Taiji, and TianQin, along with alternative orbital configurations for individual use or in network scenarios.
+As for the ground-based detectors, we examine the currently operational LVK and the third-generation detector ET, evaluating them both individually and in network configurations.
+Our analysis focuses on the performance of detectors across different polarization modes, with an emphasis on the perspective of the response function.
+Additionally, we employ the FIM to quantify ppE parameter constraints for both space- and ground-based detectors by simulating three typical masses of MBHB and SBBH.
+We present the results under various combinations and further explore how multiband and multimessenger observations enhance ppE constraints, providing a comprehensive analysis from diverse perspectives.
+For space-based detectors, Taiji provides the most stringent constraints on GW polarization modes, with LISA outperforming TianQin.
+Specifically, TianQin exhibits significantly weaker constraints on scalar modes compared with LISA and Taiji, but the uncertainties in other ppE parameters are similar between the two detectors.
+In network scenarios, the combination of LISA+TJm slightly surpasses other combinations in performance, and a network comprising all three detectors outperforms networks consisting of only two detectors.
+For ground-based detectors, the differences in constraints on frequency-related ppE parameters are negligible, while LIGO’s results for vector modes outperform those of Virgo and KAGRA, approaching the performance level of ET.
+In network scenarios, the LVK combination does not surpass the results of ET alone, with the LVK+ET network providing the most optimal constraints..
+Overall, ground-based detectors lack discrimination between breathing and longitudinal modes and exhibit significantly lower capability in constraining scalar modes compared with space detectors.
+Apart from this distinction, the constraints on ppE parameters from both space- and ground-based detectors are comparable.
+For multiband observation, it can effectively resolves the degeneracies encountered by ground-based detectors in breathing and longitudinal modes, addressing their inability to discern scalar modes.
+This approach also enhances the capability of Virgo and KAGRA to constrain vector modes, reaching a level comparable to LIGO and offering certain improvements for other ppE parameters.
+Furthermore, TianQin exhibits the most significant enhancement, followed by Taiji, with LISA demonstrating comparatively lesser improvement.
 For multimessenger observation, it provides the most significant enhancement for TianQin, bringing its results for vector modes close to those of LISA.
-For scalar modes, LISA and Taiji see greater improvements.
-ET shows a larger improvement compared to LIGO, and Virgo and KAGRA experience substantial enhancements, with their results for vector modes approaching the level of LIGO.
-Overall, both multiband and multimessenger observations can enhance GW detection results to a certain extent, leading to better constraints on GW polarization modes.
-In future research, we plan to approach the study from several aspects.
-First, we will use longer-duration GW signals, which can contain more information, and consider the non-light-speed propagation caused by different modes, where vector and scalar modes may arrive before or simultaneously with tensor modes [70].
-Second, we aim to delve into the effects brought by Time-Delay Interferometry (TDI) technology, as different TDI combinations may result in subtle differences in the outcomes [71].
-Additionally, we will consider various sources, examining the impact of ppE parameters on unequal mass BBH systems.
-Lastly, in the context of multimessenger observations, we will study the performance of EM detectors under realistic conditions, providing results that are more aligned with actual scenarios to enhance GW observations.
-In summary, through these in-depth studies, we aim to advance the research on GW polarization modes, evaluate detector performance more comprehensively and in greater detail, and provide more valuable results.
+Regarding scalar modes, both LISA and Taiji witness greater improvements.
+ET demonstrates a larger improvement compared with LIGO, while Virgo and KAGRA experience substantial enhancements with their results for vector modes approaching the level of LIGO.
+Overall, both multiband and multimessenger observations can augment GW detection outcomes to a certain extent, leading to improved constraints on GW polarization modes.
+The future research will be approached from multiple perspectives.
+Firstly, we will utilize longer-duration GW signals, which possess a greater capacity for information content.
+Additionally, we will consider the propagation of GWs at speeds other than that of light due to different modes.
+It is worth noting that vector and scalar modes may arrive before or simultaneously with tensor modes [77].
+Secondly, our objective is to delve into the effects brought about by time-delay interferometry (TDI) technology.
+This entails exploring the subtle differences in outcomes resulting from various TDI combinations [78].
+Moreover, we will examine diverse sources and assess the impact of ppE parameters on BBH systems with unequal masses.
+Thirdly, our plan involves employing specific modified theories of gravity in conjunction with existing GW events to yield more stringent constraint results [79, 80, 81, 82, 83, 84].
+Lastly, within the context of multimessenger observations, we aim to investigate the performance of EM detectors under realistic conditions.
+This approach allows us to provide results that align more closely with actual scenarios and enhance GW observations.
+The aim of these in-depth studies is to advance the research on GW polarization modes, comprehensively and meticulously evaluate detector performance, and provide more valuable results.
 Acknowledgements.
 This work was supported by the National Key Research and Development Program of China (Grant No.
-2021YFC2203004), the National Natural Science Foundation of China (Grant No.
+2023YFC2206702), the National Natural Science Foundation of China (Grant No.
 12347101), and the Natural Science Foundation of Chongqing (Grant No.